### v1

# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

Abstract
Synthetic network traffic generation has emerged as a promising alternative for various data-driven applications in the networking domain. It enables the creation of synthetic data that preserves real-world characteristics while addressing key challenges such as data scarcity, privacy concerns, and purity constraints associated with real data. In this survey, we provide a comprehensive review of synthetic network traffic generation approaches, covering essential aspects such as data types, generation models, and evaluation methods. With the rapid advancements in AI and machine learning, we focus particularly on deep learning-based techniques while also providing a detailed discussion of statistical methods and their extensions, including commercially available tools. Furthermore, we highlight open challenges in this domain and discuss potential future directions for further research and development. This survey serves as a foundational resource for researchers and practitioners, offering a structured analysis of existing methods, challenges, and opportunities in synthetic network traffic generation.
Index Terms:
Network traffic synthesis, deep learning, challenges, network traffic data types.
I Introduction
With the unprecedented growth of network traffic and the variety of network-based applications in recent years, traffic models and traffic traces have become essential components for both network providers and users for tasks such as network design, capacity planning, Quality of Service (QoS) provisioning, performance tuning, and optimization. Accurate and representative traffic plays a significant role in understanding network behavior [165, 74], maintaining QoS [91, 15, 96], optimizing systems [126, 157], and benchmarking critical network systems [115]. Furthermore, network traffic analysis is crucial for securing network systems while protecting user privacy [67, 154, 131], identifying bottlenecks in real-time [6, 24], and forecasting future demands [4, 41, 114] to name a few.
While measured traffic traces and those generated using statistical models served as a main pillar of network traffic analysis in the past, the emerging design, provisioning, and management tools are based on data-driven approaches such as machine learning (ML) and artificial intelligence (AI). A major challenge for developing and training such modes is the need for large and complex data sets that represent network traffic including the interaction of myriads of technologies, protocols, and applications. For example, Deep Neural Network (DNN)-based tools typically require large volumes of network traffic data to achieve optimal performance [136, 67]. However, it is essential that network traffic data, rich in fundamental traffic characteristics, provide valuable insights into their behavior. Another example where synthetic traffic will play a critical role is Digital Twin technology, which is envisioned to be an enabler for innovation in 5G/6G networking in many scenarios including industrial IoT and V2X systems  [101, 84]. In many circumstances, creating the necessary conditions on physical networks for gathering appropriate data for such systems may not even be possible. For example, a digital twin may be developed prior to the actual deployment of a novel 6G configuration, or one may need to create multiple disruptions to evaluate the robustness of a configuration. Such cases can be handled only via synthetic traffic traces. Even with a dedicated network, collecting high-quality traffic from real-world settings remains challenging.
First, the sheer volume of data transmitted through networks makes it difficult to extract relevant traffic while filtering out noisy background data. Second, deploying data collection tools in complex network environments often requires extended periods, leading to delays in completing network traffic analysis tasks [36]. Third, controlling the traffic sources to operate a network under the necessary conditions for trace gathering or for creating anomalies may interfere with normal operation or even be impossible to achieve. Additionally, stringent data regulations and privacy policies further complicate the process. These regulations often restrict operators and researchers from freely collecting, sharing, or analyzing traffic data, as doing so could expose sensitive information [119, 54].
To address these challenges, researchers have increasingly turned to synthetic data generation. The use of synthetic data to meet data challenges is not unique to networking, as can be seen in its success in fields such as medical image-based diagnosis [180, 29], object detection on autonomous vehicles [72] and fault detection [160, 20]. This approach has been proven instrumental in achieving optimal and robust outcomes in network traffic analysis. Synthetic network traffic generation frameworks are designed to create artificial datasets that preserve the characteristics of real-world data while mitigating issues such as data privacy, security concerns, and noise in the data. By leveraging synthetic data, either independently or in combination with limited real-world data, many network traffic analysis tasks (such as aforementioned) have demonstrated enhanced performance [137, 136, 68].
Reviewing the recent research literature, we observe that a plethora of work has been done in this domain, which we identify under two primary categories. The first category includes early methods based on statistical models, which analyze the statistical characteristics of network traffic features and generate new data that conform to these properties. The second category consists of generative AI models, such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Diffusion Models (DMs), and transformer-based techniques. These methods, supported by deep neural networks (DNNs), have significantly advanced synthetic traffic generation by capturing complex network traffic patterns from training data and replicating them in synthetic traces.
Additionally, we identify and review work that extends these two approaches to network traffic simulators, which often integrate statistical or ML/AI tools to generate network packet streams while preserving protocol-specific attributes. Furthermore, some work has been extended to commercially available network traffic generators that incorporate configurable parameters and are applied in both research and industrial settings.
To highlight the growing interest in the field of network traffic generation, Fig. 1 presents the number of publications in this domain from 2000 to 2024, categorized into four key approaches: statistical methods, Deep Learning (DL) techniques, network simulation, and commercially available generators. The articles were retrieved through a Scopus search using relevant keywords in the titles, abstracts, and keyword sections. Notably, since 2018, there has been a marked shift toward DL-based methods, as evidenced by a sharp increase in publication rates over the past five years.
Figure 1: Research papers in network traffic generation
Motivated by these findings and the growing interest in the field, this survey paper provides a detailed examination of advancements in network traffic generation. To our knowledge, it is the first survey to systematically explore and compare a wide range of synthetic network traffic generation techniques, including traditional statistical methods, DL-based models, network traffic simulators, and commercially available traffic generators. This analysis provides a broader view compared to earlier surveys, which mainly focused on specific generative models like GANs [3, 118]. For example, we begin by reviewing the statistical approaches, followed by Machine Learning-based models, like Autoregression and K-nearest neighbor clustering. Next, we examine DL-based methods, giving more attention to this area due to the growing interest and their ability to handle complex traffic patterns including those for time-varying networks. These include Autoencoders, GANs, DMs, and Transformer-based architectures. We also look at different data formats and several open challenges in network traffic generation followed by future directions.
To summarize, our contributions are as follows:
  • Comprehensive review across different methods: We provide a detailed analysis of various network traffic generation approaches, spanning from statistical methods and DL-based models to extensions into network traffic simulators and commercially available network traffic generators. Through this, we provide a broad picture of how these different methods have evolved in terms of model architectures being used in data generation.
  • Analysis of different data types used: We provide a detailed taxonomy of different data types of interest for network traffic generation. This analysis is particularly based on what information resides in data samples and how the real data is collected at different network layers.
  • Identification of open challenges: We also identify several open challenges, particularly regarding aspects such as privacy concerns, computational efficiency and complex protocol standards, etc. This discussion provides guidelines for further growth in research in this field.
  • Exploration of future research directions: We discuss the potential commercialization of advanced generative models, along with the need for synthetic data techniques to handle unseen network scenarios for enhanced system resilience.
The remainder of the paper is organized as follows: existing surveys related to this research area are reviewed in Section II. Section III discusses various data types used in network traffic generation. The review continues with an exploration of statistical methods in Section IV, followed by an in-depth analysis of deep learning-based approaches in Section V. We present simulations and commercially available tools in Section VI. Open challenges in network traffic generation are addressed in Section VII. Finally, in Section VIII, we provide hints for future research directions, and concluding remarks are provided in Section IX.
II Related Surveys
Previous surveys have largely centered around the use of GANs for generating network traffic data across a wide range of applications, with GANs long being considered state-of-the-art (SOTA) for such tasks. For example, the authors in [23] reviewed how GANs have been used for network traffic generation, categorizing network traffic under time-series patterns and distinguishing it from other types such as flow and sequence generation. This work highlights GANs’ effectiveness in simulating network behaviors with time-dependent characteristics. Bourou et al. [21] reviewed the use of GANs for tabular data synthesis, particularly for generating intrusion detection system (IDS) datasets, where GANs were shown to model the behavior of network attacks effectively. Similarly,  [9] explored the evolution of GANs, demonstrating their role in generating network flows and showcasing them as critical tools for future developments in both network traffic generation and the broader field of cybersecurity systems.
Building on these insights,  [3] assessed the application of GANs in creating synthetic attack data, emphasizing their contribution to building more robust cybersecurity defenses. This line of work positions GANs as essential for generating high-quality data that can be used for stress-testing security systems. The authors in  [56] took a broader approach by conducting a systematic review of generative ML models, including GANs, autoencoders, and hybrid models, and investigated how these models could be leveraged for intrusion detection. Moreover, the work in [118] conducted a comparative study focusing on different GAN architectures, exploring their effectiveness in generating malicious cyber-attack data. The study identified the strengths and limitations of different GAN variants in replicating diverse attack behaviors.
In parallel,  [2] provided a comprehensive survey of network traffic generators, categorizing them based on their functional behaviors and classifying how researchers have used these tools across different network simulation scenarios. This broader focus, while still heavily centered on GANs, also touched upon other data generators, highlighting their growing role in network traffic analysis and performance testing.
In contrast to existing surveys on network traffic generation, which have primarily focused on specific model types, particularly the GANs, this review broadens the scope by analyzing a diverse range of generative models, including statistical methods, DL approaches, and simulation-based platforms. In particular, we cover recent advancements with SOTA generative models, such as DMs and Transformers, which have not been addressed in previous surveys. Additionally, it examines critical aspects such as data types, evaluation metrics, and open challenges in the field. Furthermore, we provide insights into how the network traffic generation domain is evolving and advancing toward the future.
III Data types in network traffic generation
Figure 2: Different network traffic data types/formats used and data collection points in network traffic generation are categorized according to the network OSI layer.
Figure 3: Extraction of payload, packet level, flow level, and raw signal using a .pcap. The sample .pcap is from [159]
In this section, we provide a detailed overview of different data formats used in network traffic generation. We further study different data collection tools used in network vantage points. Fig. 2 shows the overview of this categorization according to different network OSI layer and taking a sample packet structure to indicate different information extract at each layer.
III-A Different data types or formats
We identify three main data formats in network data generation namely: Payload, Packet header data, Raw signal and Network usage conditions.
III-A1 Payload
In payload generation, actual byte values related to the message carried in the network packet (i.e., added at the Application layer) is generated. These byte values are typically represented as a word (2 bytes) in hexadecimal notation on tools such as Wireshark [159], Tcpdump [151] as shown in Fig. 3. In literature we observe that generation of payload has got less attention compared to other data types (e.g., packet header data) due to two main reasons. First, byte value patterns in payload data is highly complex with higher variance. Therefore, training a generative AI model that can learn these complex patterns is significantly challenging. Even if such patterns are generated, they cannot be used for reconstructing original application data. Second, for many network related downstream tasks, having packet header information is sufficient to get accurate analytics outcomes. For example, any changes in the payload are replicated in packet length parameter in the packet header. Note that we have not found major related work that cover HTTP/HTTPS data generation that mainly comes under Application layer protocols.
Despite the aforementioned challenges, the paper, PACGAN [30] proposes a GAN-based payload generation approach. In this approach, authors first separate the two digits of the hexadecimal notation of a byte. Then, each digit is mapped to a sub-range of serial values (i.e., one-to-many mapping) which will be fed to the GAN model for as the input. The output sequences are re-mapped to the original hexadecimal notations to create the byte values. Each generated byte value is then duplicated to $d\times d$ blocks forming a large matrix by concatenating each block. The final byte values are derived taking the mean of the blocks.
In [55], Guo et al. pre-process the data by removing protocol headers and extracting the first 784 bytes from the application layer. Packets exceeding 784 bytes are truncated, while those smaller are padded with 0x00 values. The authors argue that focusing on the payload retains the packet’s original information, and limiting it to the first 784 bytes ensures real-time performance. A similar method to generate payload is proposed in [156], however, the authors have fixed the length of the byte vector to 1480 bytes.
We also observe that the simulation based network traffic generators including NS3 [7], iperf [129], ositano [121], TRex [33], and Scapy [147], also create data packets with payloads. While some tools, such as Scapy, enable users to customize payloads by entering specific messages, others generate payloads randomly, simply following the rules of different network protocol. This is because, the main purpose of these simulators is to test the network performance, for example, to measure key metrics such as bandwidth, latency, and jitter and test the speed and reliability of network links, optimize configurations, etc.
III-A2 Header based generation
This is the most common network packet component used for data generation. Observing literature, we identify three main reasons why header-based generation is important compared to other data formats (e.g., payload, bit sequences from physical layers, etc.). First, packet header contains summary statistics of the network packet. For example, packet length or packet size, protocol information, and packet arrival time are some of those key information that we can easily extract from the header. Second, this header information is useful and sufficient for many downstream tasks, particularly in ML-driven analysis of network traffic data. For example, ML-based network traffic classification models to identify the usage of different applications [94, 64], user activities [39] etc., take these packet header information as the input. Third, with the availability of range public (e.g., Wireshark,  [159], Tcpdump [151]) and proprietary tools (e.g., SolarWinds Network Performance Monitor [141], Cisco Meraki  [34]), extraction and monitoring of unencrypted packet header information is easy.
Information extracted at different layers:  As shown in Fig. 2, there are three main network OSI layers used for packet header extraction and used for network traffic generation. While the packet header from each layer contains various parameters, we highlight only the mostly used key information for data generation below.
  1. 1.
    Transport layer: Commonly known as TCP/UDP header or datagram header, Transport layer packet header contains source port, destination port, Flags (SYN, ACK, FIN, and RST) [64, 171].
  2. 2.
    Network layer: Also known as segment header. The total length of the packet, source and destination IP addresses, time-to-live (TTL) and protocol field in the upper layer (e.g., TCP or UDP) are the commonly used Network layer information for data generation [171, 167].
  3. 3.
    Data link layer: Source and destination MAC (Medium Access Control) addresses, ether type (i.e., protocol encapsulated in the upper layer) and frame length are the common features extracted [40].
While several works [137, 136] have focused on generating data extracted from a single layer (e.g., frame length from the Data Link layer), other research has also explored generating data from multiple layers. For example, in  [171, 64], the authors generate multivariate data by extracting source and destination IP addresses and protocol information from the Network layer, along with source and destination port data from the Transport layer.
Table I: Comparison of selected properties of packet and flow level data and their impact on data generation process
Property Packet level Flow level Impact on data generation
Granularity High granularity offer more details of packet data e.g., packet header, payload. Low granularity as only a summary of the selected packet stream is given. Depending on the granularity requirement, the complexity of the generation task can vary. For example, packet levels with higher granularity may require complex model architectures to learn patterns with higher variability.
Amount of data High amount of data as each packet is considered Low amount of data as a summary of each flow is recorded [61] High-volume packet-level analysis captures detailed variations but may slow model convergence [61], while lower-variability flow-level data enables faster convergence with smaller datasets [89]
Memory usage Requires more memory to handle large amount of data Comparably low memory requirements [61] During model training packet level data needs more memory with large batch sizes and training time.
Processing overhead More processing requirements to process a large number of packets Less processing requirements with smaller datasets For more processing requirements additional resources including CPUs, GPUs may be required.
Purity of data With more packets, higher possibility of having noisy packet data Impact of noisy data is often diluted by data aggregation Noisy inputs will be reflected in the output data reducing the quality of synthetic data [164]
Table II: Data collection tools at OSI layers of the network
OSI layer Tools used Basic details
Application Layer HTTP Proxies (e.g., Burp Suite [120], Fiddler [149]) Act as a proxy which captures or intercepts HTTP/HTTPS traffic. Different capturing modes are provided to capture local host or remote host traffic within the same local area network. [150]
Wireshark [159] Provides encrypted version of payload added at the application layer in hexadecimal notation.
Transport Layer Wireshark [159] Provides network packet captures related with various transport layer protocols (e.g., TCP, UDP, SCTP) with easy-to-handle Graphical User Interface (GUI). Can work in different operating systems.
Tcpdump [151] A lightweight Command Line Interface (CLI) tool for network packet capture that collects packets under the same protocols from the transport layer. Mostly supports Unix-like system.
Network Firewalls (Cisco ASA (Adaptive Security Appliance) [145, 146]) Sit between the internal network and external network (e.g., internet). Monitor and control packets particularly for detecting anomaly and malicious behavior.
Network Layer Router and Switch Monitoring (e.g., ManageEngine OpManager [95]) Especially designed for network layer device monitoring. Conduct port level analysis, device failures, and create switch/router level logs.
Wireshark and Tcpdump Both tools capture network layer protocols including IPv4, IPv6, ICMP, ARP etc.
Network IDS (e.g., Snort [139], Suricata [143]) Detect malicious packet streams with rule-based mechanisms while enabling packet capture, statistical analysis, and profiling.
Data Link Layer Wireshark and Tcpdump Both tools capture data link layer protocols including Ethernet, WiFi, etc.
Network TAPs (Test Access Points) [38, 142, 144] Physical device which can capture Layer 2 packet data. Copies the traffic through a given link to a monitoring device.
Physical Layer Network TAPs Can capture the raw electrical signal bit-by-bit before any further processing. Capturing happens in a non-intrusive manner without interfering with the signal flow
Software Defined Radios (SDR) (e.g.,RTL-SDR [112], HackRF One [53]) SDR contains its basic components as in software implementation replacing the traditional hardware components. This adds more flexibility in capturing and analyzing wireless signals from various technologies: WiFi, radio, Bluetooth, ZigBee, etc.
Packet and Flow level data: As shown in Fig. 2, header-based information is used to create mainly two data formats: Packet and Flow level data. In packet-level data, the raw network packet capture, usually in .pcap file format, is used for data generation [87, 45, 57, 64, 30]. As shown in Fig. 3, a given .pcap file is a time-series sequence of network packets and each record (or raw) of the packet capture contains multiple features from different layers described before in this section. Once a data generation model is trained with packet capture data, the model is expected to synthesize packet captures with the same properties and these properties can vary significantly based on user application, network bandwidth conditions, connection medium used [69, 87].
In flow level data, we create summary statistics of selected network packet stream (see Fig. 3). This packet stream, a.k.a. flow, is defined by a tuple of five elements: source and destination IP addresses, source and destination port addresses, and protocol [82]. In addition to these five elements, a given flow often contains summary statistics of the packet stream including the total number of packets, total bytes, average throughput, etc. in both uplink and downlink directions [35]. Generally, flow-level data generation models take a collection of such flow-level records as the input and synthesize similar records maintaining the same flow-level properties. Similar to packet-level data, flow-level properties vary based on similar conditions we highlighted in the previous paragraph.
Packet and flow-level data have their own advantages and disadvantages in terms of network traffic generation. We compare and contrast selected properties of these two data types and how those affect the data generation in Table I. For a better comparison, we assume packet-level and flow-level generated using the same packet capture.
III-A3 Signal generation
Data transmission signals are present in the Physical layer, where data is sent through either wired (e.g., Ethernet, fiber optics) or wireless (e.g., WiFi, radio waves, Bluetooth) channels. In these mediums, signals are transmitted as electrical or optic pulses and wireless waves. Compared to all other upper-layer data, generating the patterns in the physical layer is significantly challenging because of the high variability and no semantic meaning in the signals.
Despite these challenges, the generation of physical layer signal is attempted in [169, 111, 113, 134]. The authors in [111, 113] present a physical layer modulation and a coding scheme taking codeword indices or bits from radio signal as the input. They leverage conditional GANs for this purpose which compares two conditional probability distributions representing ground truth and approximated distributions. In [134], GANs are employed to generate adversarial signals capable of initiating spoofing attacks during wireless data transmission. The adversary intercepts these signals and constructs ground truth datasets by positioning a malicious receiver close to the legitimate receiver, enabling it to capture signals with nearly identical channel characteristics.
III-A4 Network usage conditions
As illustrated in Fig. 2, network usage conditions include various conditional attributes that influence network traffic transmission. For instance, at the Application layer, factors such as the type of application (e.g., web browsing or video streaming) or user behavior (e.g., whether usage occurs during the day or night) impact the overall nature of transmitted network traffic. For example, while video streaming patterns contain periodic chunks of bytes downloading, IoT traffic may show non-periodic pulses sending short data streams to the IoT servers [137]. At the Network and Transport layers, the available bandwidth plays a key role in determining network throughput. Therefore, generative models that incorporate these metadata elements can enhance the contextual relevance of synthetic data.
It is important to note that a given attribute may serve as a condition for one network traffic trace but not for another [87]. For example, the packet time (i.e., indicating the time of the day or peak and off-peak time) might influence available network bandwidth, thereby affecting synthetic traffic patterns as a conditional factor. However, this same information could also function as standard packet header information generated by the model which will be used in header-based data generation. Thus, conditional attributes are highly context-dependent and vary according to the data generation scenario.
Very little work incorporates explicit conditional network traffic generation. In [87], conditional data (i.e., introduced as metadata by the authors) is collected as a ground truth component alongside related network traffic data. For example, in one of their datasets, which includes network packet-level features such as round-trip times and packet loss rates, factors like internet connection technology, ISP, and client location are considered metadata. They utilize an auxiliary GAN model to generate this metadata, which is then used as conditional input to generate the main network traffic data.
III-B Data collection devices/vantage points
Collecting a sufficient amount of network traffic data is important to have high-quality synthetic data. This can be achieved by deploying appropriate data collection tools at the required network layer. Table II reports a basic analysis of such data collection tools at different network layers. Note that several tools such as Wireshark, Tcpdump, and network TAPs provide network data information at different network OSI layers. For example, Wireshark provides packet content from the Application layer (encrypted payload) to the Data link layer. While we have found HTTP proxy tools that can collect HTTP/HTTPS, to the best of our knowledge no related work could be found in HTTP data generation.
IV Statistical methods based generation
Figure 4: Overall workflow of major statistical methods used.
In the early stages of synthetic data generation, statistical methods emerged as a foundational approach to modeling, predicting, and synthesizing network traffic. Research efforts focused on understanding the underlying characteristics of network traffic distributions to identify patterns that define network behavior. Fig. 4 provides a high-level overview of the workflow for using statistical methods for data generation.
First, captured network traffic samples are processed to extract statistical properties of traffic, such as arrival rates, inter-arrival times, packet sizes, etc., and based on the characteristics of the extracted features, a suitable distribution/model (such as Poisson process, Markov Model, etc. that are discussed in this section) is chosen. Next, a model/distribution will be generated by estimating the corresponding parameters to fit the captured data, and the generated distribution/model will be validated against the real traffic. Once a suitable model is established, it can be used to predict future traffic patterns by changing the model parameters as necessary. These models were used to predict future traffic behavior and inform network planning and performance analysis. Additionally, synthetic traffic traces could be generated based on these models to simulate various scenarios and evaluate network performance.
Next, we present the literature found for synthetic data generation using statistical methods, categorized by each statistical model type. Initially, the fundamental principles of each model are presented, followed by an analysis of how these models have been applied to network traffic generation across various application scenarios. Table III provides a summary of these studies, detailing the application scenario, the statistical model used and its adaptations, the type of data analyzed, and the availability of associated artifacts.
Table III: Summary of works leveraging statistical methods.
Work Application scenario Method(s) used Data Artifacts
OSI layer Format/Type available
Paxson et al. [117] Modeling wide area traffic Poisson Model Transport Session/ Flow/ Packet No
Ammar et al. [7] Generating realistic IP traffic Poisson Pareto Burst Process Network Flow Yes
Yoonsen et al. [172] Generating data center network traffic Poisson Short-Noise Model Network Flow No
Arfeen et al. [11] Modeling traffic in Internet access and backbone core networks Weibull Distribution Transport/ Network Session/ Flow/ Packet No
Lee et al. [79] Modeling TCP/IP traffic of wireless local area networks Piecewise Weibull Distribution with Canonical Multinomial Cascade Model Transport/ Network Flow/Packet No
Arfeen et al. [12] Modeling Internet traffic Weibull Distribution (two-parameter) Transport/ Network Session/ Flow/ Packet No
Muscariello et al. [105] Modelling Internet traffic 5-parameter Hierarchical Markov Modulated Poisson Process Transport/ Network Session/ Flow/ Packet No
Xie et al. [163] Modeling time-varying oscillation behavior of Internet traffic Nested Hidden Markov Models with Variable State-duration Transport/ Network Arrival rate process No
Dainotti et al. [43] Source-based modeling of internet traffic st packet level Hidden Markov models Application Packet No
Garcia et al. [50] Detecting malicious network behavior Markov Chains Transport Flow No
Shu et al. [135] Predict network traffic FARIMA Network Packet No
Zhou et al. [179] Predicting network traffic ARIMA/GARCH Transport Packet No
Pashko et al. [116] Modeling of self-similar traffic and loading process of telecommunication networks Fractional Brownian Motion Not specified Not specified No
Meng et al. [98] Modeling control plane traffic of 5G networks Two-level Hierarchical State-machine Application Control plane events No
Cieslak et al. [32] Handling imbalance in IDS K-nearest Neighbors Not specified Not specified No
Vu et al. [153] Handling data imbalance for encrypted traffic identification K-nearest Neighbors Network Flow No
IV-A Poisson process
This is a mathematical model used to describe a sequence of random events that occur independently and at a constant average rate over time or space. It is widely used in fields like queuing theory, telecommunications, and physics to model random phenomena such as arrivals, failures, or requests. The key characteristics of a poison process includes independence (occurrences of events are independent of each other) and stationarity (the probability of an event occurring in a given interval depends only on the length of the interval, not its position in time or space). This process can be defined by:
  • A rate parameter $\lambda>0$, which represents the average number of events per unit time or space.
  • The number of events $N(t)$ in a time interval $[0,t]$, which follows a Poisson distribution as follows where $\lambda t$ is the expected number of events in the interval $[0,t]$:
    $P(N(t)=k)=\frac{(\lambda t)^{k}e^{-\lambda t}}{k!},\quad k=0,1,2,\ldots$
Early research in modeling network traffic and synthesizing focused on using the Poisson process to accurately model the dynamics of network traffic. A major factor for utilizing the Poisson process, rooted in the principles of statistical independence, proved effective for certain scenarios [85]. Research was done to explore the limitations of Poisson modeling and identified scenarios where it failed to accurately represent the observed traffic patterns [117]. In the early 1990s, two groundbreaking papers revealed that both Wide Area Network (WAN) and Local Area Network (LAN) traffic exhibit Long Range Dependence (LRD) and self-similarity [80], [117]. In addition, the authors of [117] show that Poisson processes are valid only for modeling the arrival of user sessions (TIELWET connections, FTP control connections), but they perform poorly for other WAN arrival processes. Therefore, it became evident that the Poisson process was inadequate for this task. Recognizing the limitations of the Poisson process, researchers initially explored modified approaches based on the Poisson process to capture LRD and self-similarity properties.
One such alternative model used is the Poison Pareto Burst Process (PPBP). Ammar et al. [7] introduced a new tool for generating realistic internet traffic named NS-3, an open-source discrete-event network simulator for internet systems. By utilizing the Poisson Pareto Burst Process, the paper focused on capturing the LRD characteristics observed in real-world traffic. The traffic model was validated using an autocorrelation function plot that revealed the LRD property of real-world internet traffic.
The limitations of the Poisson process were more evident when modeling high-speed network traffic, particularly real-time data traffic modeling for next-generation networks. For example, Liji et al. [85] demonstrated that the Stationary Poison Increment Process can only model Short Range Dependence (SRD) but not LRD. To address this limitation, the authors proposed using second-order self-similarity models, such as fractional Gaussian noise and fractional ARIMA processes, as a more appropriate approach.
In the meantime, researchers also explored modeling data center network traffic using poisson processes. To better simulate realistic traffic in data center environments, the generation of flow-level network traffic matrices based on the poisson shot-noise model is proposed in [172]. By incorporating factors such as flow arrival rates, intra-rack traffic ratios, flow sizes and durations, the poisson shot-noise process offers a more accurate representation of traffic patterns in data centers.
IV-B Weibull distribution
As discussed earlier, the limitations of Poisson processes for modeling network traffic led to exploring other distributions. One such promising model was the Weibull distribution, mainly due to its flexibility to model both heavy and non-heavy tailed distributions [11]. The Weibull distribution is a continuous probability distribution where the probability density function is as follows:
$f(x)=\frac{\gamma}{\alpha}\left(\frac{x-\mu}{\alpha}\right)^{\gamma-1}\exp\left(-\left(\frac{x-\mu}{\alpha}\right)^{\gamma}\right)$
Here, $x\geq\mu$ and $\gamma,\alpha>0$ where $\gamma$ is the shape parameter, $\mu$ is the location parameter and $\alpha$ is the scale parameter. The case where $\mu=0$ and $\alpha=1$ is called the standard Weibull distribution, while the case where $\mu=0$ is called the 2-parameter Weibull distribution.
Initial studies focused on the application of Weibull distributions to wireless TCP/IP traffic. Kullback et al. [79], highlighted the inadequacy of conventional distributions in modeling the marginal distribution of wireless TCP flow inter-arrival times. The authors proposed a piecewise Weibull distribution and a canonical multinomial cascade model to capture the salient characteristics of wireless TCP flow inter-arrivals. This model achieved an acceptable accuracy for predicting packet loss rates. However, it states that while wireless TCP/IP can be partially modeled, particularly the inter-arrival times of new flows and the number of packets per flow, the arrival process of packets within a flow is not tractable due to correlations, periodicity, and inconsistency.
Further research was done in modeling internet traffic across various network tiers utilizing the Weibull distribution. Arfeen et al [12] focused on how the two-parameter Weibull distribution can be used in internet traffic modeling. It demonstrates how the Weibull distribution captures the transformation of inter-arrival process of all structural components of internet traffic (packets, flows, and sessions) as traffic moves from access to the core network. Furthermore, the authors show that given a suitable scale parameter (media or tier specific) the Weibull shape parameter can be used to zoom in from session to flow and to packet-level inter-arrivals.
Later in 2019, A. Arfeen et al. [11] provide the analytical justification for why the traffic converges to Weibull distribution as traffic moves from access to core links. This convergence is attributed to the Weibull distribution’s flexibility in capturing the stochastic properties of internet traffic at different levels (packets, flows, sessions). The study presents a comprehensive analysis of all structural components of internet traffic, linking fractal renewal processes to the modeling of correlated traffic. Additionally, the research validates Weibull-based count and interval traffic data modeling using both data fitness and queuing analysis.
IV-C Markov Model
This is a mathematical framework for modeling systems that undergo transitions from one state to another in a probabilistic manner based on the Markov property. Markov property states that the future state of the system depends only on the current state and not on the sequence of events that preceded it.
In contrast to the previous models, Markov models offer a complementary approach, particularly capturing the temporal dependencies and oscillatory behavior often observed in network traffic. Despite the inability to achieve true LRD or self-similarity, Markov models can be used to model internet traffic due to the possibility of exploiting analytical techniques to predict network performance, which according to [105] is important as it is the ultimate goal when adopting models to either study existing networks or design new ones.
Most work that proposed various traffic models to capture LRD prior to the work of L. Muscariell et al. [105] lacked practical application in network design, planning, and performance analysis has been limited due to the complex mathematical structures underlying these stochastic processes. The authors further highlight that long-term correlations beyond a certain threshold may not significantly impact system performance and hence suggest simpler models like the Markov Modulated Poisson Process (MMPP) can be effectively employed in many cases. Accordingly, the authors introduce a 5-parameter hierarchical MMPP to model real-world internet traffic in multiple levels, including sessions, flows, and packets. Three of these parameters can be directly mapped onto average traffic parameters, such as the average flow arrival rate, the average number of packets per flow, and the average arrival rate of packets within flows. The other two parameters, the average number of flows per session, and the arrival rate of new sessions define the notion of session and are used to control the Hurst parameter of the synthetic traffic on the considered scaling range. The authors conclude that the behavior of this synthetic traffic in a queue with either finite or infinite buffers is coherent with real-world internet traffic.
In 2012, Y. Xie et al.  [163] proposed a novel approach to model the oscillatory behavior of network traffic using Nested Hidden Markov Models (NHMMs) with variable state-duration. The proposed model synthesizes traffic traces that match the input trace’s characteristics such as temporal structure of oscillating behavior, statistical properties, self-similarity, queuing behavior, and multiscale energy distribution. The authors address the limitations of traditional Hidden Markov Models (HMM) and MMPPs, which can suffer from overfitting and increased computational complexity when increasing the number of states. NHMMs, on the other hand, provide a more effective and flexible approach. The model employs a hierarchical structure, consisting of two nested hidden Markov chains and an observable process. The first-layer hidden Markov chain, with variable state-duration, controls the overall oscillatory process. The second-layer hidden Markov chain models the local fluctuations within the oscillatory process, conditioned on the first-layer Markov chain. The results demonstrate the superior performance of NHMMs with variable state-duration in capturing the time structure of real network traffic. Compared to empirical statistical methods, multifractal wavelet methods, HMMs, and MMPPs, NHMMs more accurately match the oscillation characteristics, self-similarity, and queuing behavior of the input traces.
The authors in [43] utilize hidden Markov models to model inter packet time and packet size for Internet traffic from multiple application-layer protocols. The authors evaluate the model’s ability to learn, generate, and predict realistic packet-level traffic patterns through automated analysis of empirical traffic traces, focusing on the marginal distributions, auto-covariance, and cross-covariance of inter-packet time and packet size. The work shows that the proposed model is capable of predicting the short-term future behavior of the analyzed traffic based on limited monitored traffic. Unlike previous work, which primarily focused on HTTP, this study extends the application of HMMs to other application-layer protocols such as SMTP.
Among the work that presented the use of Markov chains to synthesize network traffic for practical applications,  [50] explore the application of Markov chains for detecting malicious network behavior. This analyzed the network traffic behavior between several botnets and normal network traffic and designed a behavioral network only considering the size, duration, and periodicity of each flow. By creating a dataset using virtual machines that ran real botnets, the work recognizes that the most important behaviors to be detected on a botnet are those of Command and Control (C&C) channels. All the flows of the same C&C channel are aggregated by representing a connection of flows that share a 4-tuple of source IP address, destination IP address, destination port, and protocol. The behavioral model then identifies several states for the C&C channels based on the size, duration, and periodicity for each flow and represents the transitions between states as a chain. A concatenated chain of states is then used to generate a Markov chain for the C&C channel, which is then stored on disk as a representation of that behavior. Additionally, the probability of generating the original chain of states using its own Markov chain is also stored for the purpose of botnet detection. The detection is done by finding out whether a new chain of states was generated by the Markov chain of a pre-trained model. The work recognizes that traffic from a C&C channel in one botnet will be different to other botnets. Furthermore, part of the testing chains might be very similar while other parts will be completely different. The authors designed a detection method that compares a chain of states from a 4-tuple with a Markov chain that does not have all of its states. This method obtained an F-Measure up to 93%.
IV-D Autoregression
An autoregressive (AR) model is a type of time series model where the current value of a time series is expressed as a linear combination of its previous values (lags) and a random error term. An autoregressive model of order $p$, denoted as AR(p), is defined as:
$X_{t}=c+\phi_{1}X_{t-1}+\phi_{2}X_{t-2}+\cdots+\phi_{p}X_{t-p}+\epsilon_{t}$
where, $X_{t}$ is the value of the time series at time $t$, $c$ is a constant term, $\phi_{1},\phi_{2},\ldots,\phi_{p}$ are the autoregressive coefficients, $p$ is the number of previous time steps (lags) considered in the model, and $\epsilon_{t}$ is a random error term (white noise) with mean 00 and constant variance $\sigma^{2}$. These models are widely used in time series analysis and forecasting.
ARIMA (AutoRegressive Integrated Moving Average) models extend the autoregressive model by considering the moving average and non-stationarity. FARIMA (Fractional ARIMA) extends ARIMA by introducing fractional differencing to account for long-memory dependence.
Several later works further explored ARIMA/FARIMA models and proposed improvements. Shu et al. [135] investigates the use of FARIMA models to accurately predict network traffic. The study acknowledges that for an accurate prediction, the model should be able to capture both Short Range Dependence (SRD) and Long Range Dependence (LRD) characteristics seen in network traffic. The authors state that the FARIMA models, the Markov modulated processes, TES models, and scene-based models have the capability to capture both SRD and LRD. However, the authors argue that the complexity of FARIMA model fitting has hindered their application for prediction. To address this, the authors propose a simplified FARIMA model fitting procedure. They observe that the order (p, q) of the fitted FARIMA model often remains relatively stable over time, even for extended periods. This allows for a streamlined fitting process. Additionally, the study introduces an adapted prediction method that provides an upper probability limit for the forecast. Their results highlight the suitability of FARIMA models over AR models in h-step forecasting. The study concludes that the FARIMA models offer better accuracy while requiring fewer parameters.
ARIMA/GARCH [179], a non-linear time series model that combines the linear ARIMA with conditional Generalized Auto Regressive Conditional Heteroscedasticity (GARCH) to model and predict network traffic. The authors present that an accurate network traffic prediction model should have the ability to capture SRD and LRD, Self-similarity on a large-time scale, and multifractal on a small-time scale present in network traffic. The authors identify that the exponential decay of the autocorrelation function of ARIMA captures only the SRD characteristic, but fails to capture LRD characteristics. Similarly, FARIMA models can capture both SRD and LRD but fail to capture the multifractal characteristics. Multifractal Wavelet Model (MWM), which captures multifractal characteristics, fails to predict network traffic. Furthermore, network traffic also exhibits non-stationary and non-linear properties. Addressing all the issues with previously explored models, ARIMA/GARCH model was introduced. The proposed GARCH component contributes to the dynamic variance, where variance varies over time. This conditional variance property captures the network’s bursty characteristics. Furthermore, the study presents a new procedure to estimate the parameters of the ARIMA/GARCH model and an adaptive prediction scheme. The prediction performance shows that this model outperforms the traditional FARIMA model.
IV-E Fractional Brownian Motion
A fractional Brownian motion (FBM) is defined as a centered Gaussian process $B_{H}(t)$, $t\geq 0$ with the covariance function,
$\mathbb{E}[B_{H}(t_{1})B_{H}(t_{2})]=\frac{1}{2}\left(t_{1}^{2H}+t_{2}^{2H}-|t_{1}-t_{2}|^{2H}\right)$
Here, $H\in(0,1)$ is the Hurst parameter or the Hurst index.
Fractional Brownian Motion (FBM) has been considered a suitable choice for modeling network traffic due to its inherent self-similarity, a property that mirrors modern network traffic. The authors in [116] study the accuracy of simulation for network traffic using FBM. The authors describe simulation methods for FBM based on its spectral representation and the property of stationary increments. The indicated methods were used for the simulation of self-similar traffic and the loading process of telecommunication networks. The simulated FBM is then used as an input process to a linear system. The goal is to construct a model that accurately approximates FBM while also considering the system’s response. While the Gaussian nature of FBM aids in analyzing queuing behavior, its limitations in capturing short-term correlations and rich scaling properties of real traffic have led researchers to explore multifractal models [105]. However, these models were also difficult to manage due to their analytical complexity.
IV-F Other
IV-F1 Hierarchical State-Machine-Based Traffic Models
These have been introduced as a more accurate approach to modeling control-plane traffic in cellular networks, particularly for 4G LTE and 5G. Traditional models, such as Poisson processes, have proven inadequate due to their inability to capture the bursty nature and long-tailed distributions observed in control-plane traffic.
The feasibility of addressing these limitations by employing a two-level hierarchical state-machine structure to accurately model control-plane traffic was presented by J. Meng et al. [98]. The authors show that this structure enables the capture of event dependencies among different types of control-plane events. Furthermore, by using a Semi-Markov process, the model can accurately represent the time spent by a User Equipment in each state, overcoming the limitations of the Markov model. Additionally, an adaptive clustering scheme is included to account for variations in traffic behavior based on factors such as device type and time of day. The proposed model is experimentally validated by comparing the synthetic traffic generated by the model against real-world traffic traces.
IV-F2 K-Nearest Neighbors
Classifying network traffic can be a challenging task. Dataset imbalances, where categories are not evenly represented, can lead to biased classification results, particularly when using techniques that favor majority classes. [32]. Popular methods to balance the class distribution include random undersampling and oversampling techniques. However, these methods have notable drawbacks. Random undersampling can inadvertently discard valuable data points, while oversampling by replication may introduce bias and increase the risk of overfitting. [32, 153]. By considering the above issues, instead of over-sampling by replication, [27] introduced the Synthetic Minority Over-sampling Technique (SMOTE) to over-sample by using synthetic data. Synthetic samples are generated by operating on the feature space of minority samples and their k-nearest neighbors.
With the goal of improving the true positive rate of intrusions without significantly increasing the false positives on a highly imbalanced intrusion dataset, D. A. Cieslak et al. [32] evaluate and compare sampling methods including oversampling by replication, SMOTE and undersampling. Furthermore, the study introduces a clustering based implementation of SMOTE, cluster-SMOTE. cluster-SMOTE develops minority region approximations by applying k-means clustering to the set of minority examples. Then SMOTE is applied to each cluster separately. This process allows for focused improvements on a localization basis for the minority class and improves SMOTE’s performance on imbalanced datasets.
In a similar effort, [153] investigates the sampling techniques used for addressing imbalanced data by classifying encrypted and unencrypted traffic. Since the unencrypted traffic is much higher than the encrypted traffic, it results in a dataset imbalance. Apart from the earlier mentioned methods, the study also investigates the use of Condensed Nearest Neighbor for undersampling and SMOTE-SVM for oversampling.
IV-G Summary
While statistical methods provided a valuable foundation for network traffic modeling and synthesis, their limitations became increasingly apparent with the discovery of self-similarity, SRD, LRD, non-stationarity, and non-linear properties in network traffic ([117, 179]). These characteristics, along with the growing complexity of modern networks, made it increasingly difficult to accurately model network traffic using traditional statistical approaches. Moreover, the abundance of available traffic data suggested that machine learning techniques, with their ability to learn complex patterns from large datasets, might be more suitable and efficient for modeling network traffic.
V Deep learning based generation
Figure 5: Trends in research papers highlighting the focus on various deep learning models over time.
This section reviews the main deep learning models used for synthetic network traffic generation: VAEs, GANs, DMs, and Transformers. Fig. 5 shows that GANs were the most widely used model in 2022, with 24 publications. Since then, there has been growing interest in Transformers and DMs, with publications on these models increasing from 2 in 2022 to 10 in 2024. This shift indicates a move towards more advanced models like DMs and Transformers (e.g., GPT), due to their ability to generate higher-quality data and better capture the time-based patterns in network traffic.
The remainder of this section provides a summary of the studies categorized by each model type. Initially, the fundamental architecture of each model is presented, followed by an analysis of how these models have been adapted for network traffic generation in various application scenarios. Furthermore, a brief discussion is included on the challenges associated with utilizing these models for network traffic data. In addition, similar to Section IV, each study is summarized in a tabular format, reporting the application scenario, model and its adaptations, data type, and artifact availability.
V-A Variational AutoEncoder (VAE)
Table IV: Summary of VAE-based work reporting application scenario, model architecture, and the data being used.
Model Work Application Scenario Model Data Artifacts
Variant details OSI layer Format/Type available
Vanilla VAE Kakkavas et al.[66] Generate network traffic matrices for network management Vanilla VAE Data link Traffic metrics (total bytes/5 min) Yes
Conditional VAE (CVAE) Aceto et al.[1] Attack/Intrusion detection CVAE Transport Packet level Yes
Liu et al.[90] Attack/Intrusion detection Vanilla VAE, CVAE Transport, Network, Data link Packet level No
VAE with Sequential Models Xiao et al.[161] Modeling QoS VAE + LSTM Transport, Network Packet level traffic, QoS data (e.g., delay/loss) No
Meslet et al.[100] Generic network traffic generation VAE + RNN + GMM Transport, Data link Packet and Flow level Yes
Miscellaneous approaches Bano et al.[16] Generic network traffic generation DNN-based noise remover + LSTM + CVAE Transport, Network, Data link Packet level No
The fundamental block of VAE is autoencoders[59, 75]. Autoencoder is a neural network architecture designed to learn efficient data representations, consisting of an encoder that compresses input data into a lower-dimensional latent space and a decoder that reconstructs the original input from this compressed representation. While effective for dimensionality reduction and feature learning, traditional autoencoders lack a probabilistic framework for generating new data. The VAE, introduced in [75] in 2013, built upon autoencoders by incorporating principles from variational inference[19], which is shown in Fig. 6. In this paper, we denote this basic VAE as vanilla VAE.
Instead of encoding inputs to arbitrary points in the latent space, the VAE’s encoder outputs parameters of a probability distribution, typically Gaussian, for each latent dimension. This probabilistic encoding allows for sampling from the latent space, enabling the generation of new, unseen data. The VAE’s decoder then learns to reconstruct the input from these sampled latent representations. This architectural shift, combined with a modified loss function that balances reconstruction quality with the regularity of the latent space (using the Kullback-Leibler divergence), transformed the deterministic autoencoder into a powerful generative model capable of both compression and novel data synthesis.
In literature, we have observed that this basic VAE is often combined with other models to enhance synthetic output. For example, to maintain long-term relationships in network time-series data, RNN or its variations (e.g., LSTM) [161, 100] are integrated with VAE. Another varient of VAE is Conditional VAE (CVAE)  [16, 1], which introduces conditional labels or attributes into the model to guide the generation process. Along these variations, VAE have been used in early studies for network traffic generation. Table IV provides a summary of existing work, with detailed discussions of each study presented next.
Figure 6: Basic functionality of VAE
V-A1 Vanilla VAE
Kakkavas et al. [66] focuses on generating synthetic network traffic matrices (TM) to aid in network management and traffic engineering. TM generation is necessary because direct measurements of traffic are costly and difficult to obtain at scale. By using historical data to train a vanilla VAE, the model learns the underlying traffic patterns and generates new traffic matrices with similar characteristics, thus reducing the need for extensive real-time measurements. The VAE works by encoding input traffic matrices into a low-dimensional latent space, sampling from this latent space, and decoding the samples back into realistic traffic matrices. This approach helps network administrators synthesize traffic matrices that can be used for testing and simulations while also enabling the estimation of unobserved TMs from readily available link load data.
V-A2 Conditional VAE (CVAE)
CVAE are used for addressing the class imbalance issue, particularly in applications such as intrusion detection systems (IDS), where the adversarial action can be less prevalent compared to legitimate users. In [1], the authors provide class labels as the conditional input when training CVAE for generating packet-level data. They propose an effective binning approach of the input samples to represent the distribution of characteristics. This binning process has helped the stable CVAE training process, while reducing the impact of noise samples through aggregation of feature values at the bin level. Also, it helps preserve the privacy of the original because the reverse mapping of bins and original data is infeasible.
Similarly, the work in [90] uses CVAE to address the generation of network traffic for IDS by addressing the data imbalance problem in attack types, which negatively affects IDS performance. Taking the features extracted by CICFlowMeter [49] as the input, the VAE is employed for data augmentation, generating new synthetic samples of attack data that mimic the distribution of the original samples to increase the number of minority class samples. The CVAE, which includes class labels as the conditional input, is used to specifically generate synthetic samples for particular attack types, ensuring that each attack class is balanced in the dataset.
V-A3 VAE with sequential models
To address problems such as traffic generators not being able to adapt to all types of protocols and not generating additional information like Signal to noise ratio (SNR) authors of  [100] developed an approach named Network Clustering Sequential Traffic Generation (NeCSTGen). The authors combine VAE with RNN to generate live traffic that reproduces the original behaviour at the packet, flow, and aggregate levels. Once the VAE projects the packet features into a latent space, where packets with similar features are clustered using a Gaussian Mixture Model (GMM). The RNN is then employed to ensure coherence between consecutive packets within the same flow by learning and generating transitions between these clusters, maintaining short-term dynamics and flow-level consistency.
In [161], the authors developed LSTM-enhanced VAE to model the Quality of Service (QoS). The QoS inference problem is defined as inferring the exact value of QoS metrics at each moment given real-time traffic loads. The LSTM module is used to extract fine-grained traffic features as a condition to the VAE and the VAE module learns to reconstruct the QoS distribution from the traffic features.
V-A4 Miscellaneous approaches
In  [16], a CVAE combined with LSTM is utilized for generating synthetic network traffic and imputing missing data within traffic streams. The data generation process involves two stages. First, a DNN model is applied as a preprocessing step to remove noisy packets and retain only legitimate traffic. Second, the legitimate traffic is used to train the CVAE, with conditional inputs (e.g., traffic attributes or labels) provided at the decoder. LSTM models are employed in both the encoder and decoder to capture temporal relationships within the traffic data. The generated data aligns with the statistical distribution of the original traffic, preserving real-world dynamics even in the presence of missing input data.
V-A5 Limitations in VAE based traffic generation
Although existing research does not directly identify limitations of VAEs specific to traffic generation, their general limitations can be linked to the task of network traffic generation. First, VAEs often produce blurry reconstructions [17], which may result in the inability to replicate high-frequency components present in network traffic patterns. Second, the posterior collapse in VAE [92], where the encoder fails to effectively learn the latent space from input data, makes the decoder put greater effort into synthesizing new data. This issue affects the modeling of complex feature variations in network traffic traces, especially those with longer durations and long-term dependencies. Third, balancing the trade-off between latent space structure and reconstruction quality [86, 13] is another challenge in VAE. For instance, over-regularization may simplify complex traffic patterns, leading to less diverse and unrealistic data. Conversely, under-regularization may overly emphasize reconstruction fidelity, resulting in unstructured latent spaces and incoherent synthetic traffic samples.
V-A6 Summary
VAEs have established themselves as valuable tools in network traffic generation, effectively modeling complex network traffic distributions. Table IV provides a comprehensive summary of the work done with VAEs. From vanilla VAE to its variants, such as CVAE and combinations with RNN models (e.g., LSTM), have been applied to learn both short and long-term temporal characteristics with different network applications. These applications span quality of service prediction, anomaly detection, and intrusion prevention, covering diverse data formats, including packet and flow levels, traffic metrics, and QoS data. Despite the benefits of using VAE, challenges remain, such as posterior collapse and blurry reconstructions particularly when modeling long-term feature variations of network traffic patterns.
Table V: Summary of GAN-based work reporting application scenario, model architecture, and the data being used.
Model Work Application Scenario Model Data Artifacts
Variant details OSI layer Format/Type available
Vanilla GAN Nukavarapu et al. [109] Generic network traffic generation Vanilla GAN Transport, Network Packet level No
Cheng et al. [30] Generic network traffic generation Vanilla GAN Application Bytes in hexadecimal notation No
Noel et al. [107] Generic network traffic generation for darknet data Vanilla GAN Transport, Network Packet level No
Xiao et al. [162] Generic network traffic generation for unbalanced distributed data Vanilla GAN, Federated learning Transport, Network Packet level No
Guo et al. [55] Improve ML accuracy/class imabalance Vanilla GAN Application Bytes in hexadecimal notation No
Shahriar et al. [133] Handle class imbalance, Attack/Intrusion detection Vanilla GAN Transport, Network Packet level No
WGAN Ring et al. [128] Generic network traffic generation WGAN Transport Flow level No
Fathi et al. [48] Traffic obfuscation WGAN Transport, Network Flow level No
Zhang et al. [175] Traffic obfuscation WGAN Transport Packet level No
Mozo et al. [103] Attack/Intrusion detection WGAN Transport, Network Flow level No
Kattadige et al. [68] Improve ML accuracy/class imbalance WGAN Transport Packet level Yes
Madarasingha et al. [94] Improve ML accuracy/class imbalance WGAN Transport Packet level Yes
Sivaroopan et al.[137] Improve ML accuracy/class imbalance WGAN Transport, Network Packet level Yes
Roy et al.[130] Generic RF signal generation WGAN Physical Raw signal No
CGAN Jiang et al. [63] Generic network traffic generation CGAN Transport, Network Flow level No
Wang et al. [156] Improve ML accuracy/class imbalance CGAN Application Bytes in hexadecimal notation No
Kotal et al. [78] Improve ML accuracy/class imbalance CGAN Transport, Network Flow level No
Almasre et al. [5] IoT traffic generation for cybersecurity applications CGAN Transport, Network Packet and Flow level Yes
Sequential Model Variants Lin et al. [87] Handle class imbalance / Privacy-preserved data sharing GAN, LSTM Transport Flow level Yes
Yin et al. [171] Handle class imbalance / Privacy-preserved data sharing GAN, LSTM Transport, Network Packet and Flow level Yes
Zhang et al. [173] Improve ML accuracy/class imbalance GAN, Transformer Transport, Network Flow level No
Shahid et al. [132] Attack/Intrusion detection Autoencoder, GAN, LSTM Transport Packet level No
Qu et al.[123] Attack/Intrusion detection GAN, LSTM Transport Packet level No
Misc. Approaches Anande et al. [8] Generic network traffic generation Vanilla GAN, CTGAN, Copula GAN Network, Transport Flow level No
Meddahi et al. [97] Generic network traffic generation for SIP protocol DCGAN Application Bytes in hexadecimal notation Yes
Yang et al. [168] Generate network traffic Swing GAN, Autoencoders Transport, Network Flow level No
Kim et al. [73] Simulation tool development GAN-based tool Transport, Network, Data-link Packet level No
Dowoo et al.[45] Handle data shortage in cyber-domain Graph GAN, Image GAN, Sequence GAN, Makrov models Transport, Network Packet level No
Khan et al.[70] Handle class imbalance BEGAN N/A N/A No
V-B Generative Adversarial Networks (GAN)
GANs, introduced by Goodfellow et al. [52] in 2014, represented a groundbreaking approach in generative modeling. As shown in Fig. 7, GANs consist of two competing neural networks: a generator and a discriminator. The generator’s objective is to capture the data distribution and generate realistic samples, while the discriminator’s role is to differentiate between real samples from the training data and fake samples produced by the generator. The two models are trained simultaneously in anadversarial process. The generator attempts to produce data that closely resembles the real data, aiming to fool the discriminator. Meanwhile, the discriminator works to improve its accuracy in distinguishing between genuine data and the data generated by the generator. This adversarial setup results in a continuous improvement loop: as the generator becomes better at creating realistic data, the discriminator must also become more adept at distinguishing between real and fake data. The training procedure can be described as a min-max two-player game, where the generator seeks to maximize the probability the discriminator makes a mistake, while the discriminator attempts to minimize the mistakes.
Although GANs were initially used to generate images, researchers began applying them to time-series data. This shift has enabled the generation of network traffic, as the structure and characteristics of network traffic closely align with those of time-series data. We identify five main GAN-based model architectures in network traffic generation literature.
Figure 7: Basic functionality of a GAN
V-B1 Vanilla GAN
The basic GAN model, which comprises of only generator and discriminator modules, has shown promising results in maintaining network traffic properties, despite the sequential nature they present, in network traffic generation. The authors in [109] focused on creating a protocol-agnostic framework for packet generation developing a platform named MiragePkt. In their approach, network packets were modeled as sequences of bytes. To create these sequences they leverage packet tokenization techniques, and GANs were employed to understand the conditional distribution of bytes within these sequences. To properly understand the properties of sequence, MiragePkt consists of 1D CNN layers. They further apply post-processing techniques to convert the output byte data from the GAN model to a hex stream.
Generating bytes in hexadecimal form through basic GAN has been studied by Cheng et al.[30]. They devise an encoding scheme to convert network traffic data into matrix representations suitable for use with CNNs, which serve as the internal layers of the vanilla GAN model. This encoding technique involves transforming packet bytes (in hex format) into a matrix format (image), facilitating the integration of GANs with CNN architectures. Deviating from the conventional network traffic data, Noel et al. [107] proposed a method for generating darknet-related network traffic, crucial for the early detection of malicious activity. The approach utilizes a vanilla GAN combined with a feature selection mechanism to pre-process and clean the data effectively.
Advancing the usage of vanilla GAN models, Xiao et al.[162] proposed a federated learning-based framework for distributed GAN-based network traffic generation with self-supervised learning. The approach involves local model training at the edge, where each edge node contains multiple generators, a discriminator, and a classifier. Each generator is trained on a local dataset, while the discriminator helps improve the performance of all generators at the edge. Locally trained discriminators and generators from all edge nodes are then aggregated globally using the FedAVG algorithm. The classifier at each edge provides pseudo-labels, enabling self-supervised training of the GAN models and facilitating autonomous training of the entire system.
A fundamental purpose of network traffic generation is to address data imbalance for ML-related applications. In [55], the authors utilize a basic GAN setup to address data imbalance among various protocols and applications, noting that simple oversampling of minority classes does not yield satisfactory classification accuracy. They first pre-train a classification model with the initial imbalanced dataset and then refine it using synthetic data generated by GANs, improving the performance of classification models. They leverage bytes collected at the Application layer in hexadecimal form for each packet as the data for generation, by vectorizing and removing the packet header data from network packets.
In other network applications, vanilla GANs are utilized to generate network traffic data to enhance network intrusion detection, particularly when class imbalance is prevalent in anomalous data. In this effort, Shahriar et al. [133] apply an encoding approach to convert categorical input features into a low-dimensional continuous space. They also employ feature normalization techniques and reduce feature dimensionality using Principal Component Analysis (PCA) to support the GAN model training process.
V-B2 Wasserstein GAN (WGAN)
WGAN is a GAN variant that incorporates the Wasserstein-1 loss (i.e., earth mover distance) into its training loss function, offering more stable training and addressing the vanishing gradient problem commonly observed in vanilla GAN models. In generic traffic generation using WGAN, Ring et al. propose a flow-level data generation approach to address the challenge of handling both continuous and categorical data in network flows [128]. Before training the WGAN model, they employ three techniques to convert categorical variables into a continuous domain: treating categorical attributes like IP addresses and ports as numerical values, creating binary attributes from categorical features, and using the IP2Vec [127] method to learn meaningful vector representations of categorical features. These transformed features are then used to train the WGAN model, employing the two-time-scale update rule proposed by [58].
In  [48], the authors leverage WGAN to obfuscate network traffic, reducing the detection capabilities of various applications. Initially, they derive statistical features for the flow level data they collected followed by normalization across the flows. Then, they use WGAN to collect network traffic data and generate flow-level statistics, followed by applying WGAN again to synthesize network packet traces based on these synthetic flow-level statistics to produce obfuscated traffic. The model includes multiple WGAN modules trained on different types of network traffic based on applications (e.g., Google Chrome, Google Drive), enabling application-level traffic obfuscation.
To achieve traffic obfuscation through a camouflage approach, [175] enhance the WGAN model by integrating a self-attention mechanism, enabling it to better capture feature patterns. The camouflaged features are designed so that the modified TCP packet streams, after embedding these features, meet predefined thresholds for both the Jaccard index and cosine similarity when compared to the original packet streams.
WGANs are also used in [103] for network traffic anomaly detection. Here, WGANs are used to synthesize flow-level data aimed at detecting anomalous traffic related to cryptomining attacks, where attackers exploit computational resources without consent. To mitigate the mode collapse issue, they implemented adaptive mini-batches and introduced noise into the discriminator to prevent it from converging too quickly. This approach allowed the generator more time to improve during training. Additionally, they incorporated a multi-point single-class embedding strategy for sampling noise, utilizing uniformly distributed centroids within the noise space.
In [68], features are extracted from packet-level network traffic data. During pre-processing, network packets are grouped into predefined intervals (bins), and statistical features, such as total bytes downloaded or uploaded, are derived. Percentile mapping is then applied to scale the feature values to a range of 0–100%, reducing their variability and enhancing the training process of the WGAN model used for data generation While this approach helps structure the data for better synthesis, it processes the original traces individually, leading to increased computational overhead and longer processing times. To mitigate this computational overhead, the same authors propose effective algorithms to control the WGAN model training in [94]. This algorithm compares the fidelity between the original and synthesized data at pre-defined epoch intervals and stops the model training once the data fidelity achieves the expected levels.
Researchers have applied the idea of converting network traffic data into the image domain, where WGANs show great promise, after making multiple modifications to improve GANs for 1D data generation. In [137], the authors employ the Gramian Angular Summation Field (GASF) imaging mechanism to transform certain attributes of network traces into visual representations [158]. These images are utilized for network traffic generation through a simple WGAN framework. This imaging approach effectively captures the correlations between time series points, which previous studies attempt to model using more complex GAN architectures incorporating LSTMs and other networks. The authors develop a post-processing mechanism for the generated images and evaluate the model across multiple traffic types, demonstrating improved performance with the straightforward WGAN for image generation.
The use of WGANs also extends to the generation of radio frequency (RF) signals. However, training GANs on long time-domain sequences of RF signals presents significant difficulties. To address this,  [130] proposed representing RF signals in the spectral domain, allowing the use of existing GAN architectures. The authors consider 2.5s long 4G LTE signal samples at 40 MSamples/sec which were transformed into 256x256 compact spectrograms, where the horizontal axis represented the time dimension over approximately 2 ms, and the vertical axis represented the frequency offset within a 40 MHz band. Each spectrogram served as input to the GAN, effectively framing the RF signal generation as a 2D problem similar to image synthesis. The authors also experimented with direct training of GANs on time series data. They found that while this approach could generate raw I/Q samples, it was significantly more challenging to capture the fine-grained features and multiscale details of LTE signals compared to the spectral domain representation.
V-B3 Conditional GAN (CGAN)
CGANs extend the original GAN framework by incorporating additional information (called conditions) into both the generator and discriminator, enabling more controlled and targeted data generation. Several papers have taken this advantages in network traffic generation. For example, Jiang et al.[63] propose an IoT traffic generation approach for flow-level data in Non-Terrestrial Networks, providing traffic class (network application label) as the conditional input to the CGAN. During pre-processing steps, IP2Vec is used to transform categorical variables (e.g., IP addresses) into an embedding space, which is combined with pre-processed statistical features from the flow. For the synthesized categorical data, the nearest neighbor from the real data is identified based on cosine similarity between the embedding spaces of the synthetic and real data.
Preserving class-specific details in synthetic data is essential for achieving higher classification accuracy in downstream machine learning tasks. CGANs address this by incorporating conditional inputs into the generation process. In [156], the authors use class labels, encoded as vectors, as conditional inputs for the CGAN model. During pre-processing, packet headers are removed, and the remaining data in bytes is normalized. The resulting time series is then transformed into an image format before CGAN training. Their evaluations demonstrate that the CGAN-based approach achieves superior results compared to vanilla GAN methods.
Kotal et al. [78] introduced a domain knowledge-guided CGAN-based approach for network traffic generation, addressing the challenge of limited real data in GAN training. Knowledge-based rules are incorporated as conditions to the CGAN in learning subtle feature variations and requirements (e.g., specific port numbers for different protocols such as port 80 for HTTP), reducing the need for additional data for CGAN training. The method leverages the concept of a unified cybersecurity ontology to develop a knowledge base tailored to network traffic data.
Almasre et al. [5] utilize CGANs to generate IoT network traffic, including various types of cyberattacks, to improve the analysis of security vulnerabilities. Their approach involves pre-processing steps to encode categorical features, scale numerical features, and select the most relevant attributes for generation. The CGAN model uses ground truth labels combined with an input noise vector to generate realistic and diverse network traffic. In their evaluation across various attack types (e.g., SYN Flood, UDP Flood, Reconnaissance), they demonstrate that the proposed method achieves high data fidelity, particularly in scenarios with a very limited number of network attack samples.
V-B4 GANs with sequential models
Network traffic data, such as packet-level data, is inherently a time-series distribution. Using DL models that capture sequential properties (e.g., RNNs, LSTMs) significantly enhances the learning of time-series features. Lin et al.[87] use LSTM as generators within GANs to effectively capture correlations between points in the feature time series. Their approach includes min-max normalization to prevent mode collapse. The model also generates metadata as an auxiliary output, which serves as a conditional attribute for generating real feature values. In a later study, the same authors enhance this method by synthesizing packet and flow header data across multiple measurement epochs in [171]. This improvement addresses the limitations of tabular-based GAN methods that fail to capture inter-epoch correlations. Their method generates both flow and packet-level data, with packet-level generation handled separately for each flow. Building on the approach proposed by [87], the authors in [155] adapt the model for network traffic generation in the context of malicious traffic detection. They train the model with both malicious and benign traffic to generate both types and analyze their distributions effectively.
Transformer models are also integrated with GAN models for improved performance. Zhang et al. [173] use Relative GAN (Relativistic GAN) along with transformer modules. The Relative GANs enhance the discriminator by predicting the probability that a given sample is more realistic than other samples, rather than directly classifying a sample as real or fake. In this approach, the generator is not only responsible for producing realistic samples but also for maintaining the relative authenticity differences between samples. The relativistic discriminator improves training by reducing mode collapse and promoting diversity in generated outputs. Additionally, the internal structure of their model incorporates transformer modules to effectively capture temporal relationships, further enhancing the generation of time-dependent data, such as network traffic.
The authors in [132] explored the integration of autoencoders and LSTMs with GANs to enhance generation quality, particularly for improving NIDS. In their framework, an autoencoder was first trained to learn a latent representation of real sequences of packet sizes. This autoencoder consists of LSTM layers to capture temporal dependencies in sequential data, particularly the ordering of the packet sizes in a sequence. Subsequently, GAN was trained on this latent space to generate latent vectors, which could then be decoded into realistic sequences of packet sizes. In a similar approach,  [123] propose a GAN model combined with a LSTM approach for intrusion detection in fog computing network environments. After collecting network data, noise is removed, and features are filtered before generating data using the GAN model. This processed data is then used to train an LSTM model for detecting anomalous traffic in fog computing networks.
V-B5 Miscellaneous approaches
In this section, we review the literature that explores the use of other GAN variants and the integration of different model architectures. In [8], the authors use Conditional Tabular GAN (CTGAN) and Copula GAN for feature generation and compare their performance to Vanilla GANs both quantitatively and qualitatively. They apply different data preprocessing techniques for each GAN type: Yeo-Johnson power transform [170] for Vanilla GAN, the Reversible Data Transforms (RDT) library [102] for CTGAN, and a Gaussian Copula transformer for Copula GAN. These techniques enhance training and generation performance. Their experiments show that up to 85% of the generated (fake) data features can replace real data features without being detected, highlighting the effectiveness of their approach. Additionally, the authors in [97] employ an encoder-decoder architecture to convert packet bytes into images and use Deep Convolutional GANs (DCGANs) to generate images that enhance Session Initiation Protocol (SIP) traffic data.
The authors in [45] propose a model that combines multiple GAN architectures—Graph GAN, Image GAN, and Sequence GAN—to process different data types in parallel, generated by style-based encoders. Graph GAN processes graph data that encodes relationships between packet IP addresses. Image GAN takes image data created by converting time intervals into visual representations. Sequence GAN handles packet-layer information (e.g., Ethernet, IP) as numerical sequences. Additionally, they generate optional data, such as packet loads, using Sequence GAN and Markov models. The synthetic data produced by these trained models is decoded to recreate .pcap files.
Yang et al. [168] focus on generating network traffic using traffic obfuscation techniques. They use Gaussian Mixture Models (GMM) for data normalization and autoencoders to reduce network data to a low-dimensional form. This processed data is then fed into a GAN model that integrates Swing Transformers. Swing Transformers are a specialized type of transformer architecture designed to capture feature variations at multiple scales within deep learning layers. The use of Swing transformers has enabled the model to capture feature variations across different scales within the deep learning layers while also integrating self-attention mechanisms.
Extending their research into practical usage, [73] utilized CTGAN [166] for data generation to build a network traffic simulation platform. They began by decoding packet data to train the CTGAN model. The generated data was then encoded into synthetic packet information in the form of a .pcap stream. The quality of the generated data was assessed using a simulation setup by comparing it against real packet streams.
Khan et al.[70] proposed a data generation approach for SDN settings in healthcare consumer IoT device management, specifically addressing class imbalance. The method utilizes a Boundary Equilibrium Generative Adversarial Network (BEGAN), with a separate GAN model dedicated to each data class. The training process is halted based on a convergence metric to ensure optimal performance. Additionally, autoencoders are employed for dimensionality reduction in the data, enhancing the efficiency of the model.
In addition to the practical data generation, a thorough theoretical analysis of GAN-based data generation is studied. For example, [88] explores the privacy implications of synthetic data generated by GANs, focusing on differential privacy and membership inference attack resilience. The authors prove that generated samples inherently have weak differential privacy properties, scaling with the ratio of generated to training samples. They also show that vanilla GANs, without special privacy mechanisms, are robust against black-box membership inference attacks. Their contributions lie in providing theoretical bounds on privacy risks for GAN-generated samples, highlighting areas where traditional GANs meet privacy requirements and where further techniques are needed.
V-B6 Limitations in GAN based traffic generation
A significant challenge in GAN-based modeling is mode collapse, where the generator produces a limited variety of data, failing to capture the full range of variations in the training dataset [87]. This issue is particularly pronounced in network traffic generation, as traffic traces often encompass a diverse set of modes that are challenging for the generator to learn. However, this challenge can be effectively mitigated through approaches such as robust data normalization techniques and other advanced pre-processing methods [87]. Another major challenge is training instability, which arises when the generator and discriminator fail to reach equilibrium during training [13, 44]. This instability can result in synthetic traffic traces with reduced fidelity. Additionally, vanishing gradients is also a common issue in GANs when the discriminator becomes overly strong. In such cases, the generator receives weak gradient signals, hindering its ability to improve [44]. For network traffic generation, vanishing gradients can significantly affect the learning of long-term temporal dependencies in network traffic traces and lead to insufficient generalization of the data, as smaller gradients prevent the generator from learning comprehensive feature patterns.
V-B7 Summary
GANs are among the most widely used generative AI models for network traffic generation. Table V provides a comprehensive summary of the work done with GANs. To optimize these models for traffic generation, various GAN variants have been proposed, particularly by modifying their internal neural network architectures. These adaptations include integrating autoencoders, recurrent neural networks (RNNs), and incorporating conditional information. Such enhancements have successfully facilitated traffic generation for a wide range of network-related applications, including intrusion detection and improving other machine learning tasks, primarily for both packet-level and flow-level data. While GANs inherently face challenges such as mode collapse and training instability, existing approaches have effectively addressed these issues through advanced mechanisms, enabling the generation of high-fidelity data.
V-C Diffusion Models (DM)
Table VI: Summary of DM-based work reporting application scenario, model architecture, and the data being used.
Model Work Application Scenario Model Data Artifacts
Variant details OSI layer Format/Type available
DDPM Sivaroopan et al. [136] Anomaly detection, real-time classification, traffic fingerprinting DDPM Transport, Network Packet level Yes
Chi et al. [31] Generic RF signal generation Hierarchical Diffusion Transformer Physical Raw signal data Yes
Stable DM Jiang et al. [64, 65] Generic network traffic generation Stable DM + ControlNet [178] Transport, Network Packet level Yes
Controlled DM Zhang et al. [176] Mobile app usage modeling and service interaction tracing Hierarchical DM + Word2Vec + Encoder-Decoder + Two-layer transformer Transport, Application Network flow traces, App usage sequences No
Chai et al. [25] Network optimization DM + Attention-based classifier-free guidance + Contrastive learning Application Aggregated mobile traffic data across geographic areas Yes
Chai et al. [26] Mobile traffic generation, Network optimization Conditional DM + Knowledge graph + Frequency attention mechanism Application Mobile traffic time-series, Urban data (POI, population, regional functions) No
Qi et al.[122] 5G network traffic generation CANDLE + GCN Application Traffic time-series, Geospatial grid-based traffic patterns No
Jasni et al. [62] Adversarial packet generation, Evasion of NIDS DM + BERT Transport, Network Flow level No
Lightweight DM Li et al. [83] Synthesis of malicious traffic for edge devices Lightweight DM with depthwise separable convolutions Network Packet level No
DMs were first introduced to the deep learning community in 2015 by [140], where they were initially conceptualized as probabilistic generative models. However, it was not until 2020, with the groundbreaking work in [60], that DMs gained significant traction. The development and training of these models in this period marked a turning point in generative modeling, with DMs quickly emerging as a compelling alternative to long-dominant GANs.
DMs operate based on a probabilistic framework that transforms random noise into meaningful data samples. As shown in Fig. 8, this is achieved through a two-step process. First, the forward process gradually adds Gaussian noise into the input sample ($x_{0}$) until $T$ time steps, creating a series of noisy data representations from ($x_{1}$ to $x_{T}$). The model is then trained to reverse this process, step by step, by learning the conditional probabilities that describe how to transform a noisy version of the data back into a less noisy version. The DNN model, a.k.a. denoising model, which conducts this transformation, learns the underlying noise probability distribution of the data. Compared to VAE and GAN, the iterative denoising process in DM provides more stability during the model training while offering a diverse representation of the synthetic data, mitigating issues such as mode collapse. Different extensions of DM have been proposed so far for network traffic generation as we have summarized in Table VI and explained next.
Figure 8: Basic functionality of DM
V-C1 Denoising Diffusion Probabilistic Model (DDPM)
DDPMs optimize the denoising process by explicitly modeling the noise distribution, allowing for sharper and more diverse samples. DMs were first adapted for network traffic generation in [136], where the authors employed DDPM to model time series data representing packet-level features, specifically downloaded byte streams. In addition to time series data, they also utilized GASF images, which were derived from the same time series data. A comparison was drawn with the SOTA GANs that had previously been used for network traffic generation, focusing on both 1D features and 2D images. Their results demonstrated a significant improvement in the accuracy of network traffic image generation using DMs, outperforming previous GAN-based methods, particularly in cases that leveraged packet-level features in 2D image format.
Generating high-fidelity RF signals presents unique challenges due to the need to capture their full structure across the time, frequency, and complex value domains. In  [31], the authors highlight that RF signals, as time-series data, capture dynamic details that standard diffusion models struggle to synthesize. Moreover, critical RF features embedded in the frequency domain, such as Doppler shifts, are often overlooked due to the focus of DMs on spatial and amplitude information. Furthermore, the complex-valued nature of RF signals, including phase data, is neglected in existing models. Addressing these complexities, the authors introduce RF-Diffusion to synthesize RF signals utilizing a novel Time-Frequency Diffusion (TFD) theory and a Hierarchical Diffusion Transformer (HDT). In TFD theory, they present RF-Diffusion model which alternates between adding noise in the time domain and blurring the frequency domain during the forward process, and reverses these effects by denoising and deblurring. This restoration emphasizes the accuracy of the amplitude in the time domain and the continuity of the frequency domain. The HDT incorporates a hierarchical architecture to decouple spatio-temporal dimensions, attention-based diffusion blocks with enhanced Transformers for feature extraction, and a complex-valued design to encode both amplitude and phase information. The authors synthesize Wi-Fi and Frequency-Modulated Continuous Wave (FMCW) signals using RF-Diffusion which outperforms typical DDPM, DCGAN, and CVAE-based methods.
V-C2 Stable DM
Stable DM generates controlled high-quality images from textual descriptions, enabling both text-to-image and image-to-image transformations. Further, Jiang et al. [64, 65] applied stable DMs to address the problem of network data scarcity. The authors converted network packet traces into images through the use of nPrint, a method that standardizes bit-level information from packet headers into a consistent format. These images were then utilized with a stable diffusion model to synthesize network traffic, offering a controlled synthesis process. A key innovation in this work was the integration of ControlNet[178], which provided an additional layer of control, ensuring that essential network characteristics, such as the number of bits per header, were preserved in the synthesized images. To further enhance the realism of the generated data, a post-processing mechanism was introduced to maintain adherence to protocol-specific constraints and characteristics.
V-C3 Controlled DM
Expanding the scope of diffusion models in network traffic generation and cybersecurity, recent research has focused on improving controllability, reducing data dependency, and capturing complex spatial-temporal patterns - key areas where traditional GAN-based approaches often fall short. Zhang et al.  [176] presented a hierarchical, service-guided, controlled diffusion model for network flow trace generation. This approach was designed to explicitly model users’ dynamic network usage patterns, focusing on the interactions between users and network services, such as mobile application usage traces. Their hierarchical generation structure consisted of multiple layers, the first of which was responsible for modeling app usage behavior. This app usage data was then used to guide the generation of network flow traces, ensuring that the generated flows reflected realistic service usage patterns. The model captured co-usage relationships between different services, as well as sequential usage patterns, by leveraging a pre-trained embedding model and an encoder-decoder architecture. Furthermore, it incorporated a two-layer transformer network to model the temporal and feature correlations inherent in multidimensional network flow data. The generation process followed a two-step structure: first, a diffusion model generated the app usage patterns, and subsequently, a second diffusion model was trained on network flow traces, using the app usage patterns to guide the generation of realistic network traffic data.
In [25], the authors introduce OpenDiff, a model that leverages publicly available data such as satellite imagery, POI distributions, and population density to generate mobile traffic data for network planning and optimization. By adopting a multi-positive contrastive learning algorithm and an attention-based classifier-free guidance module, OpenDiff mitigates the reliance on non-public data and enhances the controllability of traffic generation. Data pre-processing steps included feature extraction from satellite images and POI, population distribution analysis, and multi-positive sample construction for contrastive learning.
The paper, CANDLE [122], uses a conditional diffusion model to simulate 5G network traffic in regions with limited historical data, conditioning on existing 4G traffic patterns. This model incorporates graph convolutional networks (GCN) and a cross-attention mechanism to address data scarcity and provide operators with insights for base station deployment. In addition, in [26], KG-Diff is proposed to incorporate urban contextual features from an Urban Knowledge Graph (UKG) to guide mobile traffic generation. By employing a frequency attention mechanism that decomposes network traffic into multiple periodic components, KG-Diff effectively models complex temporal patterns and the underlying correlations between urban environments and network traffic, offering a more nuanced approach to mobile traffic generation. Fourier transform was used for frequency decomposition.
Extending the application of DMs to cybersecurity, DiffuPac [62] innovatively combines a Bidirectional Encoder Representations of Transformers (BERT) model with a diffusion framework to generate adversarial network packets capable of evading detection by NIDS. DiffuPac introduces a targeted noising mechanism and classifier-free guidance, ensuring that only malicious packet sequences are noised while preserving the integrity of normal traffic. Additionally, the model employs a unique packet concatenation strategy to seamlessly integrate malicious packets into benign traffic flows, leveraging contextual embeddings from BERT for enhanced mimicry. DiffuPac’s fine-tuning phase optimizes a variational lower bound that prioritizes the reconstruction of malicious packets as indistinguishable from normal traffic patterns. Evaluations demonstrate that DiffuPac not only outperforms traditional approaches like GANs and LSTMs but also operates under the realistic constraint of limited attacker knowledge, showcasing its practical efficacy in real-world scenarios.
V-C4 Light-weight DMs
Another angle of improvement for DMs were addressed in making the model lightweight such that it can be used for high-quality data generation at edge devices. [83] proposes a DM specifically designed for resource-constrained edge devices. By integrating depthwise separable convolutions into the diffusion process, this model achieves significant reductions in computational overhead while maintaining the diversity and quality of generated malicious traffic. The evaluations demonstrate the ability of the model to produce high-quality synthetic data, improving the training of intrusion detection classifiers even on hardware-limited edge nodes. This lightweight design underscores the potential of diffusion models to address computational challenges, enabling real-time cybersecurity applications at the edge.
V-C5 Limitations
DMs face challenges in traffic generation, including limited controllability under specific conditions and difficulties in handling multimodal data or generating realistic spatial and temporal patterns. Traditional DMs often lack mechanisms to incorporate diverse contextual features effectively, reducing fidelity and adaptability. To address these issues, frameworks like OpenDiff [25] employ classifier-free guidance and attention-based fusion to enhance controllability and integrate conditional features such as human activity data. CANDLE [122] uses cross-attention mechanisms and graph-based architectures to align 4G traffic data with 5G demand, improving adaptability in underdeveloped regions.
Additionally, DMs have high computational demands, making them unsuitable for resource-constrained edge devices [83]. Models like KG-Diff [26] address this by incorporating UKG and frequency attention mechanisms to capture urban context and multi-scale periodicities, enhancing fidelity and controllability. DiffuPac [62] integrates BERT and targeted noising strategies to generate adversarial packets that blend seamlessly into normal traffic while maintaining functionality. Lightweight designs, such as depthwise separable convolutions, reduce computational overhead, enabling deployment on edge devices. These advances improve the practicality and effectiveness of DMs in traffic synthesis and cybersecurity.
V-C6 Summary
The emergence of DMs has revolutionized the field of network traffic analysis, providing a powerful alternative to traditional generative models like GANs. Table VI provides a comprehensive summary of the work done with DMs. Innovative adaptations of DMs, such as DDPM, stable diffusion models, and controlled diffusion models, have demonstrated remarkable capabilities to generate realistic high-quality network traffic data. By effectively modeling the underlying distributions of packet-level features and user interactions with network services, DMs have consistently outperformed previous SOTA GANs in terms of accuracy and realism. The ability of DMs to capture complex temporal and feature correlations while mitigating issues such as mode collapse has solidified their position as the leading generative framework for network traffic synthesis. Further, lightweight diffusion models make the way for high-quality data synthesis in a resource-constrained environment. As research in this area continues to advance, DMs are poised to play an increasingly crucial role in enhancing network security, improving data analysis, and addressing the challenges of data scarcity in network traffic environments.
V-D Transformers
Table VII: Summary of Transformer based work reporting application scenario, model architecture and the data being used.
Work Application scenario Model(s) used Data Artifacts
OSI layer Format/Type available
Bikmukhamedov et al.[18] Generic network traffic generation, Attack detection GPT-2 Transport, Network Packet level and Flow level No
Meng et al.[99] Generic network traffic generation GPT-2 Transport, Network Packet level and Flow level No
Qu et al.[124] Network optimization and Long traffic flow classification GPT + Linear attention mechanism Network Flow level No
Kholgh et al.[71] Generic network traffic generation, IDS GPT-3 Transport, Network Packet level, Flow level No
Kong et al.[77] Cellular network traffic synthesis CPT-GPT Application Event-level sequences No
Transformers were first introduced in 2017 in [152] to address sequence transduction tasks, such as neural machine translation. These tasks involve converting an input sequence into an output sequence. The transformer model is a type of neural network that learns the context of sequential data and generates new data based on its understanding of the relationships between elements within the sequence. Transformers excel at understanding context by using a mechanism called attention, which enables the model to weigh the importance of different elements in a sequence.
As shown in Fig. 9, tokenization, embedding, encoder, and decoder (i.e., with multi-head attention and feedforward layers) are the components of a generic transformer module used for network traffic generation [152]. Taking network traffic generation as an example, this process starts by tokenizing the network traffic trace into smaller units, which are converted into vector representations through an embedding layer. The encoder processes these embeddings using multi-headed attention and feedforward layers, capturing both global and local dependencies in the traffic sequence. The decoder, which takes another segment of the network traffic trace as input, uses multi-headed attention to integrate context from the encoder and applies feedforward layers to generate the subsequent traffic sequence. This setup effectively models and reproduces the patterns in network traffic data.
Figure 9: Basic functionality of Transformers
V-D1 Generative Pretrained Transformers (GPT)
Generative Pretrained Transformers (GPT) are a specific class of transformer models initially developed for natural language processing (NLP). Their utility extended beyond NLP, finding applications in various fields, including network traffic generation. After GPT-2 was released, it was leveraged in 2020 by  [18] for synthetic network traffic generation and classification. The model treated packet size and inter-packet time intervals as two-dimensional features, mapped to a one-dimensional source space using K-Means clustering. The cluster embeddings resembled tokens used in transformers for training. By feeding the traffic class as the first token into the pre-trained model, it could generate a flow corresponding to that class.
Further advancements were made in  [99], where the authors developed NetGPT, a GPT model pre-trained on extensive network traffic data irrespective of the downstream task that needs to be performed. During fine-tuning, task-specific labels (e.g., software identification, source port generation) were added. To enhance the dataset for fine-tuning, they shuffled packet headers, as the order of headers was irrelevant for certain tasks. However, the token limit posed challenges for flow length during training and generation. To mitigate this,  [124] introduced TrafficGPT, which employed a linear attention mechanism to extend the token scope from 512 to 12,032 tokens. They also developed a reversible token representation method to enable bidirectional mapping between pcap files and tokenized representations. Metadata like MAC and IP addresses were removed to avoid overfitting.
With the introduction of GPT-3 models, authors of  [71] leveraged GPT-3 Davinci model and GPT-3 Babbage model to synthesize network traffic for various tasks. The pipeline of this process would be triggered by a user requesting network traffic either using a specific scenario (e.g., normal traffic data, a Ping-of-Death attack, etc.) or by specifying a set of network protocols (e.g., ICMP, DNS, etc.) Then, a Flow Generator, which is a Python script, produces network flow in text format, i.e. it generates the sequence and textual summary of packets that make up the network traffic, although not creating the actual packets themselves. The flow summary is then passed to the Packet Generator, which is a transformer-based model capable of creating packets using the network flow. The packet generator then creates the specified packets and writes them in .pcap format to a file. GPT-3 Davinci was used to generate training samples to train the model as it was a more accurate GPT-3 model. Then, to reduce the cost and to get faster responses while inferencing, GPT-3 Babbage model was used in the packet generator.
Similarly,  [77] presents CPT-GPT, a high-fidelity traffic generator tailored for cellular network control-plane traffic. CPT-GPT effectively addresses challenges such as stateful semantic correctness, multimodal feature relationships, and long-term data drifts by leveraging a Transformer-based architecture. Through its innovative tokenization scheme and probabilistic output mechanism, CPT-GPT generates realistic traffic while avoiding issues like state violations and mode collapse. Unlike traditional Semi-Markov models (SMMs) that require extensive manual updates to align with evolving 3GPP standards, CPT-GPT autonomously achieves comparable or superior fidelity.
V-D2 Limitations
Transformer (e.g., GPT) models face limitations in network traffic tasks due to the heterogeneous syntax of headers and payloads, especially with encrypted traffic, which complicates vocabulary construction and semantic retention. Their quadratic complexity with input length poses challenges for long packet flows, while unidirectional models struggle with bidirectional tasks like classification and attack detection. The limitations in token length and the inability to accurately reconstruct realistic traffic patterns further hinder their effectiveness [124]. To address these, NetGPT [99] employs general encoding with hexadecimal tokens, processes only the first three packets of heavy flows to reduce overhead, and improves adaptability via header field shuffling and packet segmentation. TrafficGPT [124] enhances the handling of long sequences with linear attention mechanisms (supporting up to 12,032 tokens) and reversible token representations for precise traffic reconstruction. Together, these strategies improve the utility of transformer models in various traffic understanding and generation tasks [99, 124].
V-D3 Summary
The integration of transformer models, particularly GPT, into the realm of network traffic generation represents a significant advancement in the field of artificial intelligence and data synthesis. Table VII provides a comprehensive summary of the work done with Transformers. The evolution from initial applications in natural language processing to specialized frameworks like NetGPT and TrafficGPT showcases the versatility of transformers in handling complex data structures, such as network traffic. Innovations in tokenization and attention mechanisms have addressed challenges like flow length limitations and class imbalance, leading to improved performance in synthetic traffic generation. As these models continue to evolve, they hold the promise of transforming network analysis, enhancing security measures, and providing valuable insights into network behavior through more accurate and comprehensive traffic generation methodologies.
Table VIII: Comparative analysis of synthetic network traffic generation approaches (considers only the vanilla models)
Aspect Statistical Modeling Deep Learning–Based Methods
VAEs GANs Diffusion Models Transformers
Data Requirements Moderate: tuned from real measurements, fewer data required High: requires sufficient data to learn a meaningful latent space Very High: demands large volumes to avoid mode collapse and capture diversity Very High: extensive data needed to capture fine–grained, complex details Very High: typically trained on massive corpora to learn long–term dependencies
Model Complexity Low: based on distributions and parameter estimation Moderate: encoder-decoder structure with multiple layers Moderate: requires more training, sensitive to hyperparameters and mode collapse Very High: iterative denoising with multi–step sampling increases complexity High: utilizes self–attention layers with large scale pretraining
Training Stability High: analytical models yield deterministic solutions Low: stability can be affected by the trade–off between reconstruction and regularization Low: adversarial training is inherently unstable and sensitive to hyperparameters causing mode collapse Moderate: generally stable training Moderate: with proper optimization, training can be stable
Capture Complex Patterns Low: effective for basic statistical properties but struggles with non–linear dynamics Moderate: can capture temporal features, especially when enhanced with sequential layers High: well–suited for learning complex, non–linear patterns (long–term dependencies can be challenging) Very High: excels at capturing both fine–grained and long-range temporal dependencies Very High: particularly adept at modeling long–term dependencies and sequential structures
Interpretability High: parameters and distributions have clear statistical meanings Low: latent space analysis is needed for interpretability Low: adversarial frameworks provide limited insight into internal representations Low: the iterative denoising process is less intuitive Moderate: attention mechanisms offer partial insight despite overall complexity
Computational Efficiency High: typically lightweight and fast due to closed–form calculations Moderate: inference is efficient, but training can be resource–intensive Moderate: training two networks simultaneously increases computational cost Low: iterative sampling during inference is computationally expensive Low: training large transformer models (GPT) requires significant resources
Data Fidelity Moderate: capable of reproducing basic traffic patterns, but may not capture high variability or complex interactions Moderate: produces plausible outputs but may appear over–smoothed or lack fine details High: when well–trained, generates highly realistic traffic that closely mimics real–world nuances Very High: achieves SOTA fidelity in replicating complex patterns and fine details Very High: produces coherent, context–aware sequences reflecting long–term dependencies
Reproducibility High: deterministic output under same conditions (parameters) Low: Non deterministic behavior poses challenges to reproducing results Low: As it is difficult to determine when to stop the training for better output Moderate: comparatively reproducible due to better training stability Moderate: comparatively reproducible due to better training stability
Flexibility & Scalability Moderate: tied to assumptions and less adaptable to evolving traffic characteristics High: can be easily extended (e.g. with LSTM) Moderate: conceptually flexible, though practical tuning can be challenging High: versatile in modeling diverse traffic patterns, albeit at a high computational cost High: extremely adaptable across various network scenarios and traffic types
V-E Comparative analysis of different methods
Table VIII provides a summarized comparison of the different approaches used in network traffic generation. Each aspect in the left-most column evaluates key considerations for selecting a generative model:
  • Data Requirements: Measures the amount of real-world data needed for effective training. Models with Low requirements can generate plausible traffic with minimal data, while Very High indicates a reliance on extensive datasets.
  • Model Complexity: Describes the structural and computational complexity of the model, including the number of layers, hyperparameters, and training mechanisms. Low complexity models are simpler and easier to configure, while Very High complexity models will have billions of parameters and require extensive fine-tuning.
  • Training Stability: Evaluates the ease of training a model and the likelihood of instability, such as mode collapse (GANs) or trade-offs in reconstruction (VAEs). More stable models (High) are preferred for reproducibility and reliability.
  • Capture Complex Patterns: Assesses the model’s ability to learn intricate temporal dependencies and non-linear patterns in network traffic. Low models struggle with such details, whereas Very High models excel in both short-term and long-term dependencies.
  • Interpretability: Reflects how transparent the model is in providing insights into its learned representations. Statistical models have high interpretability, whereas deep learning-based models often provide limited intuition.
  • Computational Efficiency: Considers the training and inference costs. High efficiency models are lightweight and computationally inexpensive, whereas Low efficiency models require significant resources.
  • Data Fidelity: Evaluates how closely the generated traffic resembles real-world patterns. High fidelity implies the model generates diverse and realistic traffic flows.
  • Reproducibility: Determines whether the model consistently produces the same results under the same conditions. Low reproducibility models, such as GANs, may produce varying results due to instability in training.
  • Flexibility & Scalability: Examines how well the model adapts to different network conditions and scales with increasing data. Very High scalability models can generalize across multiple network scenarios.
VI Commercially available network traffic simulators
Over the years, traffic generators have predominantly focused on evaluating network protocols and infrastructure rather than the various applications running on top of them. This has been especially true with canonical tools like the venerable Iperf [129], which is often employed to test end-to-end transport protocols, or hardware-based solutions such as IXIA[148], Spirent[37], or XENA[106] which are designed for use cases such as performing RFC 2544 compliant device tests [22]. Traditionally, the division of work between software and hardware-based solutions presented a trade-off, where software tools often fell short in their ability to reach the full core network line rate, a capability more commonly achieved by specialized hardware.
However, as networks, applications, operating systems, and COTS (commercial off-the-shelf) hardware have evolved, so too have the associated evaluation tools. In recent years, there has been a noticeable shift towards the development of both software and hardware solutions that enable the creation of flexible traffic generators at a relatively low cost. This shift has been driven in part by advancements in kernel bypassing technologies, such as DPDK (Data Plane Development Kit) and eBPF (Extended Berkeley Packet Filter), as well as the widespread availability of programmable Network Interface Cards (NICs), including FPGA (Field-Programmable Gate Array), IPU (Infrastructure Processing Unit), and smartNICs.
These technological advancements have significantly impacted the networking and distributed systems communities, leading to the development of more versatile tools. On one hand, these tools are capable of achieving performance levels comparable to those of special-purpose hardware solutions when used with commodity network cards [47, 42, 33]. On the other hand, they offer tailored traffic generation for specific workloads, such as key-value stores [177, 76, 81, 174, 10]. This dual capability has opened new possibilities for network testing, allowing for more precise and application-specific evaluations, which were previously challenging to perform with older, less flexible tools.
The landscape of traffic generation has been further enriched by comprehensive surveys such as that by Adele et al. [2], who explored the functional behaviors of traffic generators across academia and industry. Rather than focusing solely on raw performance metrics, this study categorized traffic generators into seven distinct classes, ranging from "Constant or Maximum Throughput Generators" to "Script-Driven Traffic Generators." Their analysis also spotlighted the top ten most popular or cited tools between 2006 and 2018, including widely recognized names like Iperf [129], Scapy [147], MoonGen [47], or Tcpreplay, many of which remain highly relevant today.
A significant trend observed in recent years is the evolution of legacy traffic generators from standalone tools to integral components within broader performance evaluation frameworks. This shift is especially pronounced in commercial network monitoring and evaluation software suites, where traffic generation functions as one of many tools alongside capabilities such as network tomography and periphery scanning. Examples include SolarWinds Network Performance Monitor [141], Dynatrace [46], or NetScanTools Pro [108] all of which incorporate basic traffic generation features. These implementations often rely on legacy traffic generators while integrating them with canonical networking utilities like traceroute [14] or nmap [93].
Moreover, the flexibility and performance of modern traffic generators have broadened the scope of testing possibilities. No longer confined to infrastructure-centric evaluations, these tools now support complex, high-fidelity simulations of real-world application scenarios. As network environments grow increasingly dynamic with the proliferation of cloud computing, IoT devices, and edge computing, the demand for adaptive and precise evaluation tools continues to rise.
The progression of traffic generators reflects the broader evolution of networking technologies. Modern tools, driven by advancements in software and programmable hardware, offer unprecedented performance and adaptability, enabling more nuanced and application-specific evaluations. As the networking landscape continues to evolve, these tools will undoubtedly play a pivotal role in shaping the methodologies used to ensure the reliability, scalability, and efficiency of future networks. Nonetheless, we can still see nowadays a kind of purpose specialization for each tool where for example Scapy would be used primarily in the context of debugging or proof-of-concept development whereas Spirent would be one of the tools of choice for compliance and system engineering requirement validation.
Table IX: Summary of commercial traffic generators
Work Type Protocol constraints Cost Usability Main application
IXIA [148] Hardware Datalink layer to transport layer by default, additional modules for above Expensive GUI Compliance and System Engineering
Spirent [37] Hardware Datalink layer to transport layer by default, additional modules for above Expensive GUI Compliance and System Engineering
XENA [106] Hardware Datalink layer to transport layer by default, additional modules for above Expensive GUI Compliance and System Engineering
Trex [33] Software Datalink layer to transport layer (stateless) Low/Open source GUI and script Performance and management
SolarWinds [141] Software Datalink layer to transport layer Medium GUI and script Performance and management
Dynatrace [46] Software Datalink layer to transport layer by default, need to program for more Medium GUI and script Performance and Management
NetScanTools [108] Software Datalink layer to transport layer Medium GUI Development and penetration testing
OSINATO [121] Software Datalink layer to transport layer Low/Open source GUI script Development and penetration testing
Scapy [147] Software Datalink layer to application layer Low/Open source script Development and penetration testing
In order to reflect this de-facto specialization of tools, we present in Table IX an overview of some of the most popular industry-used traffic generators. In particular, we propose to classify those tools under three broad categories: compliance and system engineering, performance evaluation and management, and development and penetration testing. As a result, we can see that due to various regulations, traffic generators for compliance and system engineering tend to be dominated by hardware-based solutions as they made sure to pass numerous certifications for a given piece of equipment. In particular, we note that nowadays, this category is still dominated by tools such as IXIA[148], Spirent[37], or XENA[106].
In the second category of traffic generator applications, the performance and network management, we note the presence of one legacy fully dedicated traffic generator application, Trex [33], and two traffic generator frameworks, SolarWinds and Dynatrace. In the case of Trex, while it started more as a testing tool, its integration with the rest of the CISCO infrastructure, such as the umbrella framework, [^1] https://umbrella.cisco.com/ made this tool use more in a performance and management application with special attention to performance analysis.
Finally, we present in the last identified traffic generator category, namely the development and penetration testing, three tools that are associated with this category. In particular, we mention in Table IX, two tools offering a modular testing capability with an easy-to-use GUI in NetScannTools and Osinato [121] frameworks. Last but not least, we mention a tool that has been gaining more attraction over the years with Scapy, where the user can easily analyze, capture, and replay any type of traffic as long as performance and high throughput are not mandatory due to its python implementation.
VII Open Challenges in Network Traffic Synthesis
While synthetic network traffic data can be a powerful tool for many applications, there remain several challenges that need to be addressed to improve its utility and effectiveness. In this section, we discuss some of the key challenges and technical limitations that complicate the synthesis process.
VII-A Privacy and Utility Trade-off
One of the most critical challenges in network traffic synthesis is ensuring privacy while maintaining the utility of the generated data. Real network traffic often contains sensitive information such as IP addresses, communication patterns, and user behavior, making privacy protection essential. However, strong privacy preservation measures, such as differential privacy, often degrade the quality of the synthetic data, leading to reduced performance in downstream tasks. This opens up to the privacy-utility trade-off.
Differential Privacy introduces noise into the data to mask individual entries, but this can distort key traffic patterns. The challenge is to find the right balance between the level of noise (controlled by the privacy parameter $\epsilon$) and the ability of the synthetic data to retain useful properties for tasks like anomaly detection or traffic analysis. Fine-tuning $\epsilon$ requires a deep understanding of the privacy risks and application requirements. They can lead to a loss of granularity in the data, making it difficult to maintain fine-grained traffic features that are essential for some applications. Some preliminary work has been done in [171, 87] to address the privacy utility trade-off by incorporating DP into the GAN generation. Authors of [104, 28] have tried to address the membership inference attacks on GANs.
VII-B Adhering to Protocol Standards
Developing synthetic network data that satisfies timestamp and protocol requirements is essential, as networks operate under strict protocol standards that govern data integrity and transmission quality. Ensuring that synthetic traffic adheres to these protocols is critical for realistic and useful synthetic data. Over 80% of the work does not consider developing network traffic generation that adheres to protocol standards. This has been addressed to a certain level in [64, 65] by integrating ControlNet into DM for data generation to adhere to protocol standards and performing a post-processing step to correct the images to ensure that they adhere to the protocol. However, there still exist certain challenges to be addressed.
VII-B1 Protocol Compliance in Data Generation
Network protocols define limits on packet sizes, acceptable payload structures, header formats, and timing between packets. For instance, protocols like TCP/IP and UDP have maximum packet size constraints and specific sequencing requirements. Synthetic data must comply with these standards to ensure usability for applications like simulation and stress testing of network systems.
VII-B2 Timestamp Synchronization and Temporal Coherence
Network data is highly time-sensitive, with applications that rely on precise timestamps and sequence orders for accurate simulations. Synthesis algorithms must handle timestamps and ensure that events in the synthetic dataset mirror the temporal coherence found in real network traffic. This challenge becomes even more pronounced in applications like intrusion detection, where real-time analysis of packet sequences is essential.
VII-C Capturing Complex Network Behaviors
Real-world network traffic exhibits highly complex behaviors, shaped by dynamic routing, load balancing, and bursty traffic flows. Capturing these nuances in synthetic data remains a significant challenge. Some efforts have addressed this to an extent. For instance, the authors of [68, 94, 137] employ packet binning to condense time series traces while preserving essential feature spaces for long traces. Similarly, Sivaroopan et al. [137, 136] leverage a GASF imaging mechanism to capture temporal correlations in packet-level features. However, despite these advancements, further challenges remain in accurately modeling the intricate characteristics of network traffic.
VII-C1 Traffic Diversity
Network traffic varies greatly depending on the type of applications, devices, and protocols in use. Synthetic data generators need to be flexible enough to capture this diversity, from low-latency traffic such as video streaming to high-throughput traffic such as large data transfers. However, existing synthesis methods often fail to replicate these variations effectively, leading to over-simplified or unrealistic synthetic traffic patterns.
VII-C2 Temporal Dynamics and Seasonality
Network traffic is not static; it changes over time due to user behavior and network conditions. Ensuring that synthetic data models these temporal dynamics, including short-term fluctuations and long-term seasonality, is a major challenge.
VII-C3 High-dimensional Feature Spaces
Network traffic data is often high-dimensional, encompassing a wide range of features such as packet sizes, inter-arrival times, protocol flags, and source-destination IP pairs. Capturing correlations and dependencies across these features in a synthetic dataset is difficult. Synthetic data may either over-simplify these relationships, lose important interactions, or become too complex to be useful for downstream tasks.
VII-D Evaluating Synthetic Data Quality
Another key challenge lies in developing standardized metrics for evaluating the quality of synthetic network traffic data. While we have discussed several metrics such as statistical similarity and performance on downstream tasks, there is no one-size-fits-all solution. There are many papers  [69, 67, 137, 136] attempting to provide high-quality synthetic data, however, there are challenges yet to be solved.
VII-D1 Lack of Benchmark Datasets
Unlike other fields such as computer vision or natural language processing, there are relatively few publicly available, standardized benchmark datasets for synthetic network traffic evaluation. This makes it difficult to compare the performance of different synthesis approaches, especially for privacy-preserving techniques.
VII-D2 Generalization vs. Overfitting
Striking the right balance between generalization and overfitting in synthetic data generation remains a major issue. Overfitting to the real traffic can lead to privacy risks, while excessive generalization may result in synthetic data that fails to capture the essential characteristics of the network traffic, reducing its utility.
VII-D3 Task-Specific Evaluation
The effectiveness of synthetic data can vary significantly depending on the specific task it is used for, such as traffic classification, anomaly detection, or intrusion prevention. This calls for task-specific evaluation metrics, but defining these metrics can be complex and may require different evaluation frameworks for each application.
VII-E Computational Complexity and Scalability
As networks expand in size and complexity, the computational demands of generating and evaluating synthetic data grow accordingly. This also poses a particular challenge for real-time network traffic generation on edge devices, where computational resources are limited. To address this, the authors of [83] introduced lightweight DMs designed for efficient data generation on edge devices. However, further challenges remain in balancing computational efficiency with the accuracy and fidelity of the generated traffic.
VII-E1 Scalability of Data Generation
Large-scale networks generate massive amounts of traffic data, and synthesizing data at this scale requires significant computational resources. Many existing synthesis methods struggle to scale effectively, leading to bottlenecks in generating synthetic data for large network environments.
VII-E2 Real-time Synthesis
In some applications, synthetic data may need to be generated in real-time to support testing and analysis under current network conditions. Achieving real-time synthesis is particularly challenging when attempting to replicate complex traffic patterns while maintaining high fidelity and privacy guarantees.
In summary, network traffic synthesis has tremendous potential but faces significant challenges. These challenges include preserving privacy without compromising data utility, capturing complex network behavior accurately, and developing robust evaluation metrics. Additionally, scalability remains an issue, as generating and evaluating synthetic data for large networks requires extensive resources. Ensuring that synthetic data aligns with protocol standards and accurately represents both typical and unseen network scenarios will further enhance the realism and applicability of synthetic traffic data for testing, simulation, and security analysis. Overcoming these hurdles is essential for the broad adoption of synthetic network traffic data across varied applications.
VIII Future Directions
As communication networks evolve to meet increasing demand and diverse operational environments, accurately modeling and replicating real-world network conditions will ensure robust performance, efficient resource allocation, and seamless user experiences. In this context, network traffic generation is expected to become crucial in optimizing network services.
One such practical application of network traffic generation can be illustrated through Ericsson’s 5G Digital Twin [110], simulated in NVIDIA Omniverse Enterprise. Ericsson employs Digital Twins (DTs) i.e virtual replicas of physical systems, to model the interaction between 5G cells and urban environments in real time. While these models can predict coverage based on the physical environment, they currently lack the ability to simulate real network conditions based on factors such as variations in traffic volume, application-specific patterns, and congestion dynamics. This is where network traffic generation becomes transformative. Creating synthetic traffic that accurately mimics real-world network behavior, enables researchers and engineers to simulate, test, and refine network performance under realistic conditions. Integrating traffic-aware simulations into 5G/6G Digital Twins will significantly enhance predictive accuracy, optimize network configurations, and improve overall service quality. Beyond twins, one may also envision the possibility of networks of synthetic traffic sources ( a digital twin without a physical counterpart) generating and inserting traffic to physical networks for tasks such as camouflaging or protecting privacy.
In the remainder of this section, we explore potential enhancements and future directions for improving network traffic generation.
VIII-A Commercialization and Multimodality of Emerging Deep Generative Models
Future commercial network traffic generators could significantly benefit from the integration of advanced generative techniques such as DMs and Transformer-based architectures (GPTs), which have set new standards in generative modeling. GPTs, being pre-trained on extensive datasets as large language models, exhibit unparalleled sequence generation capabilities. This knowledge aligns seamlessly with the requirements of time-series generation paradigms, such as network traffic data, where temporal and sequential dependencies are critical. By bridging the gap between the LLM’s word embedding space and the embedding space of time-series tokens or packet flows, GPTs can be fine-tuned to generate realistic and contextually rich network traffic patterns. Similarly, DMs, known for their iterative refinement of data from noise, offer an exceptional ability to model intricate data distributions and generate synthetic data with high fidelity. These models could be particularly effective in mimicking the statistical and temporal properties of network traffic, capturing both short-term variations and long-term dependencies.
Beyond these advancements, the future of network traffic generation lies in multimodal deep generative models that integrate multiple data types, including textual descriptions, images, structured network metadata (e.g., packet headers, flow statistics), and raw RF signals. DMs such as Stable Diffusion can be conditioned on network topology diagrams or textual descriptions of attack scenarios to generate structured protocol-compliant synthetic traffic. Similarly, transformer (e.g. Contrastive Language-Image Pretraining (CLIP) [125]) models can align textual representation of network behavior with images or other modalities, allowing for more context-aware traffic synthesis that is more interpretable and adaptive. GPTs can extend their sequence modeling capabilities beyond packet flows by incorporating visual and textual network telemetry, enabling cross-modal learning for real-time network traffic simulation. The synergy of these techniques in commercial applications could enable scalable and precise traffic generation solutions, addressing long-standing challenges such as data imbalance, privacy preservation, and real-time traffic simulation. By leveraging these SOTA multimodal generative models, the next generation of commercial traffic generators could offer enhanced capabilities, transforming how synthetic network traffic is created and utilized for cybersecurity, network optimization, and anomaly detection.
VIII-B Synthetic Data Generation for Unseen Network Scenarios
As discussed earlier, synthesizing network data that represents unseen or hypothetical scenarios is crucial for evaluating system resilience and preparedness in various environments.
Generative models can simulate unseen network conditions by learning from available datasets and extrapolating patterns for unobserved scenarios, such as extreme network loads, rare cyberattacks, or unusual traffic distributions in geographically isolated locations. Additionally, when real-world data for specific platforms or protocols is unavailable, synthetic data generation can incorporate domain knowledge to approximate these environments. However, the key challenge lies in ensuring that the generated data remains both valid and useful for downstream tasks.
One possible strategy is to follow a similar approach as Neural Style Transfer (NST) [51] which combines the content of one image with the artistic style of another. NST is based on the intuition that the content (structure and high-level features) and style (textures, patterns, and colors) of an image can be extracted using different layers of a deep convolutional network and that optimizing a new image to match the content features of one image and the style features of another generates a result that retains the structure of the content image while adopting the artistic characteristics of the style image. To follow a similar approach, we first define two attributes for each network trace, Content which refers to the actual content/payload carried by the traffic trace (such as a video stream, or website), and Domain which refers to the conditions of the network which affects the traffic. For instance, when considering a network trace for a user streaming videos, the specific video title would be the content and the domain could be the specific network the user is connected to (wired or wireless) or the specific geographical location. The notation $T_{n}(C_{x},D_{y})$ represents a traffic trace corresponding to content $x$ and domain $y$. For example, consider a scenario where traffic traces with content $C_{1}$ from domain $D_{1}$, and traffic traces with content $C_{2}$ from domain $D_{2}$ are available, but traffic traces with content $C_{2}$ from domain $D_{1}$ is not available and needs to be generated. Analogous to NST, if we can extract content-specific features from $T_{2}(C_{2},D_{2})$ and domain-specific features from $T_{1}(C_{1},D_{1})$, these features can be combined to generate a new sample $T_{3}(C_{2},D_{1})$ representing $C_{2}$ and $D_{1}$.
Unlike with images where the exact layers of a deep convolutional network to extract the content and style from images are known, to the best of our knowledge, no prior research has addressed a similar approach for network traffic traces, making this an open research problem. However, P. Smolovic et al[138] introduced the use of triplet loss to extract video-specific (content) features from network traffic traces, independent of the platform on which the video was streamed (domain). Building on this, we propose that using a similar methodology leveraging contrastive loss functions, such as triplet loss, to extract both content-specific and domain-specific features from network traffic traces is feasible. Once the content-specific and domain-specific features are extracted, the same approach as NST can be followed to generate the new sample by optimizing a random input to match the extracted content and domain.
An alternative approach would be to leverage neural networks with an encoder-decoder architecture, such as a VAE. These networks can be trained to ensure that the latent space forms distinct clusters corresponding to each domain and content type. For example, as illustrated in Figure 10, this approach can be used to create distinct clusters for each domain, and within each domain, further distinct clusters are formed based on content. Once trained, the latent vector for a traffic trace could be extrapolated in the latent space to generate a trace with a different content or domain. For example, Figure 10 shows a scenario where the model is trained on traces with content $C_{1}$ from both domains and $C_{2}$ from $D_{2}$ only. Accordingly, it will learn both $C_{1}$ and $C_{2}$ clusters in $D_{2}$ and only $C_{1}$ cluster in $D_{1}$. Once trained a trace $T(C_{1},D_{1})$ can be extrapolated in the latent space to create $C_{2}$ in $D_{1}$ to match context $C_{2}$ based on the relationship between $C_{1}$ and $C_{2}$ in $D_{2}$. Then the decoder can generate a realistic network traffic sample $T(C_{2},D_{1})$ that represents context $C_{2}$ and domain $D_{1}$.
Figure 10: Scenario-based data generation: VAE
IX Conclusion
This survey highlights the importance of synthetic network traffic generation in overcoming challenges associated with real-world data collection, such as privacy concerns, scalability issues, and the need for high-quality data. We reviewed and categorized existing approaches into statistical methods and deep learning-based models, while also examining their extensions into network traffic simulators and commercially available tools. This structured classification provides insights into the strengths, limitations, and practical applications of each approach, illustrating how the field has evolved and how these methods are being used to synthesize different scenarios.
Despite progress, there are still challenges to be addressed. Preserving user privacy, ensuring realistic data that accurately reflect complex protocol behaviors, and improving the usability of synthetic data for tasks like network security and performance optimization remain significant hurdles. Advanced AI models like transformers and diffusion models offer new possibilities for synthesizing unseen networking scenarios.
By offering a detailed taxonomy of data types, network traffic generation methods and unresolved issues, this paper aims to inspire innovative solutions and guide researchers and practitioners in advancing synthetic network traffic generation methodologies for robust and efficient network traffic analysis. As synthetic data generation continues to advance, it holds the potential to transform network traffic analysis, making it more efficient, secure, and adaptable to the ever-growing demands of modern networks.